A Complete Bibliography of Publications in Canadian Journal of Mathematics = Journal canadien de

mathématiques for the decade 1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.08

Title word cross-reference

(0, 1) [268]. (0, 2) [592]. (2, q, r) [396]. (B, N) [59]. (d − 2) [58]. (p, p) [94]. +1 [306]. −\frac{1}{2}Δ + V [438]. 1 [628]. 10 [792]. 16 [649]. 2 [17, 105, 138, 341, 166, 361, 682, 512]. 28 [593, 626]. 2^{ω} [291]. 3 [79, 383, 657, 127]. 4 [786, 316]. 8 [316]. = [395]. [−1, +1] [459]. ^{∗} [679]. ^{∗} [225]. _{10}φ_{9} [561]. A^{3} [511]. A^{4} [511]. A_{n} [474]. α [483, 598, 641]. α ≥ 0 [641]. B_{n} [474]. BhatSO_{n}⟨16⟩ [512]. BP⟨1⟩^{∗}(K(Z, 3); Z/p) [374]. C [344, 745, 75, 628, 697]. C(X, M_{2}) [633]. C[![x, y]!] [791]. C^{∗} [713, 779, 712, 789]. C^{1} [665]. C^{∗} [594, 616, 579, 537, 499, 507, 622]. C^{n} [112]. C_{0} [77, 111]. C_{n} [474]. B(H) [679]. H_{u} [82]. L^{(q,λ)} [424]. Y_{u} [82]. CM [558]. d [394]. D_{n} [474]. dim, phisbKsbXotimesL(X) = 1 [586]. E(R_{R}) [172]. E^{3} [138, 431]. F [706, 534]. F_{q}((1/t)) [711]. G [520, 762, 402]. G/K [244]. G_{2} [383]. g_{ij} [673]. Γ\G [758]. geq1 [210]. m [484]. GF(q^{3}) [566]. GL(n, q^{m}) [318]. GwrAsbn [601]. GwrSsbn [601]. H [85]. H^{1}(T) [666]. H^{2} [730]. H^{2}(L, F) [659]. H^{∞}(X) [428]. hatE′ [228]. heta [137, 331, 192]. I [427]. II [396]. II_{1} [679, 625]. III [646]. III_{0} [449]. III_{λ}; (0 < λ < 1) [449]. JC [397]. K [99, 683, 682, 665, 789, 645, 493, 297, 28]. K_{0} [779, 712]. K_{2}(O) [786]. k_{α,n} [315]. K_{m,n} [53]. Ksb0 [563]. L [118, 621, 638, 398, 168, 744]. L(f) [265, 501]. l^{1} [265]. L^{2} [730]. Ł^{p}(X) [690].

L^{p}[0, 1], 1 < p < +∞ [577]. l′ [351]. L_{1} [784]. L_{p} [224, 322, 111]. Lambda [523, 389]. Lsb2(I)oplusC^{r} [607]. lspp [500]. M [515]. M^{∗} [585]. mod [628]. N [677, 181, 478, 170, 154]. n^{2}heta; (mod, 1) [620]. n^{2} × 4 [89]. Ω [414]. Ω^{3}S^{3}X [682]. \overline{∂} [696]. p [740, 581, 612, 91, 117, 189, 771, 376, 255, 485]. P^{3} [593]. P^{3}, [626]. P^{n} [273]. pi [560, 650]. PSL(2, 31) [15]. PSL(2, R) [680]. Q [664, 503, 56, 288, 491, 777, 200, 494, 726, 541, 703, 316, 24]. R[x] [134]. R^{3} [742]. R^{k} [572]. R^{n} [66]. R^{n+1} [547]. R^{n+k} [572]. R_{ij} [673]. S [608, 22]. S(t, t + 1, v) [233]. S(X) [557]. S(Y) [557]. S^{3} [653]. s_{p}(α) [131]. scrCscrL [509]. σ [116, 239, 789]. σ_{4} [146]. SL(2, 5) [24]. SL(3) [765]. SL(r) [398]. sl_{3} [502]. SO^{∗}(2n) [419]. Sp(4) [771]. Sp_{4}(k) [605]. SPIN^{c} [691]. Ssb0(G) [611]. SU(2) [580]. τ [239]. U_{n}(V) [216]. varphi [524]. VN(G) [237]. w^{∗} [473]. x^{n} [26]. x^{n} + a [576]. Z [550]. Z_{(p)} [98].

-actions [146]. -adic [612, 189, 771, 376]. -algebra [706]. -algebras [594, 616, 713, 779, 397, 537, 499, 712, 789, 507, 622]. -analogue [777]. -analogues [288]. -analytic [28]. -ary [484]. -Azumaya [520]. -ball [166]. -Bessel [726]. -closed [85]. -closure [137]. -coherence [99]. -collapsible [341]. -colorings [170]. -compactification [509]. -connected [628, 79, 105, 361, 127]. -connectedness [116]. -construction [664]. -continuous [331, 192]. -convex [641]. -diagonals [579]. -dimensional [181, 649]. -discreteness [28]. -disfocality [483]. -embedding [75]. -endomorphisms [679]. -extension [494]. -extreme [58]. -forms [383]. -functions [697]. -group [91, 117]. -groups [740, 581, 255, 585, 712, 515, 168]. -Hermite [56]. -homology [682]. -hypercyclic [491]. -ideals [394]. -indistinguishability [398]. -inner [297]. -integral [273].

-interpolation [592]. -invariance [665]. -isolated [351]. -Jacobi [200]. -Krull [414]. -lattices [558]. -manifolds [657]. -mappings [389]. -matrices [268]. -Neumann [696]. -nodal [344]. -orderings [745]. -pairs [59]. -properties [569]. -rank [786]. -rational [703]. -regular [79]. -rings [534]. -separable [650]. -sequences [503]. -series [541, 638]. -smoothness [239]. -spaces [500, 784]. -special [560, 650]. -sphere [138]. -spiral [598]. -splines [523, 118]. -stable [402]. -stables [402]. -subgroups [22]. -théorie [493]. -theory [683, 665, 789, 621, 645, 493]. -torsionfree [172]. -valuations [745]. -versal [402]. -verselles [402]. -vertex [478].

81m [338].

A. [204, 336]. Abel [57, 554]. abelian [122, 581, 733, 241]. abéliens [662]. Absolute [75, 629, 148]. absolutely [258]. abstract [297]. action [729, 1, 124]. Actions [572, 653, 476, 746, 381, 146, 559, 743]. acyclic [136]. addition [726, 432]. additive [181, 699]. Additivity [273]. adic [612, 189, 771, 376]. admissible [446]. admit [495, 444]. admitting [238]. AF [779]. Akcoglu [69]. Albert [454]. Algebra [434, 86, 247, 471, 6, 458, 237, 150, 502, 536, 184, 706, 669, 283, 488, 721, 422, 611]. Algebraic [149, 378, 423, 699, 7, 166, 498, 586, 570, 493, 399]. Algebras [136, 536, 488, 594, 790, 740, 467, 516, 520, 616, 385, 600, 278, 543, 713, 774, 41, 618, 178, 659, 779, 397, 16, 530, 634, 350, 601, 526, 269, 140, 720, 114, 537, 750, 499, 382, 768, 220, 378, 206, 230, 389, 609, 195, 1, 124, 407, 712, 263, 789, 507, 222, 661, 313, 323, 225, 622, 477, 359, 132, 743, 454, 427, 776, 325]. Algèbres [16, 790, 407]. algébrique [493]. algorithm [48, 275]. Almost [510, 3, 61, 93, 44]. Almost-sure [44]. along [728, 764, 415]. Alternating [383, 701]. amalgams [611]. amarts [12]. Amenability [342]. amicable [203]. ample [764]. Amsterdam [338]. analogue [777]. analogues [288, 71]. Analysis [668, 553, 623, 742]. Analytic [609, 38, 152, 28, 784, 536, 442, 292, 469, 697, 225, 488]. analytical [665]. analytiques [225]. Angular [574, 298]. anguleux [298]. Antichains [562]. Any [138, 341]. aposyndetic [163]. Application [504, 384, 71, 274, 676, 748, 275]. Applications [623, 239, 784, 777, 364, 459, 494, 654]. approach [570]. approximate [589, 148, 766, 793, 359, 132]. Approximately [507]. Approximation [401, 253, 523, 628, 31, 369, 736, 584, 26, 112]. approximations [412]. AR-spaces [148]. arbitrary [152]. arcs [655]. arise [473]. arising [186]. Arithmetic [103, 711]. Arithmetical [109]. arithmétique [711]. arrays [89]. ary [484]. Askey [716]. Associated [421, 288, 280, 456, 488, 654]. associée [456]. associées [280]. asymmetric [510]. Asymptotic [362, 694, 441, 254, 517, 368, 36, 52, 453]. asymptotically [683]. automorphism [91, 257, 160]. Automorphisms [520, 656, 749, 117, 437, 604, 722]. autotopism [444]. axiom [503]. axioms [175]. axis [772]. Azumaya [520].

Baire [408, 116, 564]. ball [783, 166]. balls [138]. Banach [86, 513, 790, 553, 623, 775, 111, 153, 530, 634, 535, 692, 120, 359, 132, 460, 64]. Banach-valued [460]. band [748]. band-width [748]. base [711]. bases [234]. Basic [464, 474, 777, 541]. basis [790, 604]. basis-conjugating [604]. be [286]. Behavior [598, 254, 441, 522, 52, 519]. behaviour [624, 362]. Behrend [379]. Bergman [270, 78, 211, 302, 574].

Bernstein [141, 305]. Bessel [726]. Betti [279]. between [750, 389, 638]. Beyond [502]. Bezout [545]. bi [561]. bi-orthogonal [561]. Bianchi [671]. bidisc [730]. Bifiltered [235]. bifiltrés [235]. biinfinis [556]. bilinear [494]. bimodules [769]. binary [532]. binding [417]. bipartite [275]. Biplanar [67]. biquadratic [309, 485]. Birkhoff [292, 154]. block [233]. block-schematic [233]. Blocking [51]. blocks [286]. blow [788]. blow-ups [788]. BMO [442]. BMOA [218]. Bochner [508]. bold [600]. Boolean [385, 272, 661]. bordism [691, 735]. Borel [685, 57]. Borel-type [685]. bornage [599]. bornées [577]. Borsuk [106]. bound [657]. boundaries [553, 623]. Boundary [778, 555, 489, 700, 736, 720, 48, 415]. bounded [577, 783, 333, 171, 772, 536, 211, 68, 359, 132]. boundedly [384]. Bounds [279, 517, 169]. Boxes [66]. Braid [337]. braids [338]. branched [764]. branches [637]. branching [350]. Brauer [124, 311, 1, 29]. Brieskorn [396]. Brooks [314]. Brownian [460]. Buchsbaum [95]. Buildings [570]. bundle [19]. bundles [470].

c.n.e [299]. cables [756]. calculus [42, 591]. can [286]. cancellation [622]. Canonical [762, 422]. capacity [330]. Capitulation [94]. caps [527, 102]. cardinal [118]. Carleman [514]. Cartan [246, 447]. Cartesian [733]. case [248, 319]. categoricity [231]. categories [360, 39, 410, 192]. category [737]. Cauchy [781, 55, 580, 472]. Cech [80]. cellular [385]. center [237, 333, 132]. centers [36]. Central [359, 91, 619, 414]. centralizers [359]. Centres [525]. Certain [730, 143, 162, 785, 312, 219, 565, 473, 669, 445, 198, 498, 622, 404]. certains [198, 316]. Cesàro [685, 554, 486]. chainable [559]. chains [562, 296]. Chandra [762].

Character [650, 75, 490, 292, 404]. characteristic [88]. characterization [90, 14, 15]. Characterizations [521]. characterized [197]. characterizing [2, 123]. Characters [612, 699, 560, 746, 650]. Chebyshev [280]. Chevalley [306, 550, 213]. Choquet [720]. Chromatic [290, 307, 126]. circle [435, 523, 118]. Class [309, 268, 740, 35, 94, 236, 77, 723, 282, 414, 558, 429, 443, 311, 617, 632, 225, 130]. classe [225]. classes [104, 76, 693, 735, 443, 12, 447, 63, 739]. classical [454]. classical-Albert [454]. Classification [586, 648, 743, 208, 597, 323]. Classifying [789]. Clifford [296]. close [770, 617]. close-to-convex [770, 617]. Closed [220, 87, 553, 623, 232, 37, 448, 85, 462, 506, 661, 404]. closure [476, 373, 137]. Coabsolutes [190]. codes [162, 8, 627]. codimension [17, 139]. Coefficient [4, 375, 104]. Coefficients [312, 366, 115, 688, 598, 196, 641]. Cohen [526, 97]. coherence [99]. coherent [324]. Cohn [204, 336]. cohomology [676, 717, 512]. collapsible [341]. collineation [495, 649]. Collineations [482]. colorings [170, 221, 65]. Colour [606]. coloured [165]. column [646]. column-finite [646]. combinatorial [60, 63]. comma [627]. comma-free [627]. Commutative [324, 533, 164, 656]. Commutativity [686, 242]. Commutator [667]. commutators [464]. commuting [683]. Compact [178, 649, 370, 463, 476, 384, 473, 457, 673, 261, 201, 574, 559, 743, 317, 427]. compact-range [457]. compacta [431]. compactification [80, 155, 509, 331]. compactifications [260, 339]. Compactness [782, 457, 514, 385]. complementary [356]. Complete [6, 434, 445, 217]. Completely [163, 313, 481, 763]. completion [151, 405]. completions [55, 525]. Complex [332, 788, 219, 751, 441, 530, 634]. complexes [103, 539]. complexity [747]. composés [198]. Composition [783, 437, 574]. compounds [198]. concave [450]. concave-convex [450]. Concerning [116, 395, 685, 546]. condition [284, 42, 591, 549]. conditions [143, 476, 277, 265, 75, 676, 534, 242, 228]. cone [13, 185, 395, 742]. cones [110, 411, 712, 704]. configurations [167, 296]. Conformal [62]. Congruence [119, 543, 90]. Congruences [763, 672, 377, 587, 368, 161]. conjecture [435, 106, 170, 593, 626, 29]. Conjugacy [337]. conjugate [89]. conjugated [301]. Conjugates [707]. conjugating [604]. conjuguée [301]. connected [628, 79, 105, 261, 361, 127]. connectedness [116]. connectivity [165]. conormal [19]. constant [104, 596]. constrained [575]. construct [93]. constructible [317]. constructing [275]. construction [259, 664]. contain [111, 329]. contained [329]. containing [536, 689]. context [670]. context-free [670]. continua [684, 462, 506, 163, 559]. continuation [697]. Continuity [548]. Continuous [87, 33, 232, 778, 55, 736, 509, 596, 331, 68, 750, 197, 258, 200, 561, 747, 192]. contractibility [504]. contractions [500]. Contractive [633]. control [42, 700]. Convergence [708, 12, 701, 662, 153, 480, 400, 154]. convex [675, 435, 510, 433, 478, 40, 473, 450, 535, 304, 770, 617, 411, 120, 661, 468, 641]. Convexity [534, 5, 64, 532]. convexoid [145]. convolution [618]. convolutions [453]. Coprime [746]. Corner [519]. corona [776]. corps [352, 198, 316]. Correction [124, 626, 506]. correspondence [676]. correspondences [243, 650]. Corrigenda [61]. Corrigendum [471, 123]. countable [75]. counterexample [122]. Counting [165]. covariant [721]. Cover [92]. covered [286]. covering [268, 156]. coverings [764, 391]. covers [327]. Coxeter [306]. Criteria [239, 187, 485, 108, 688, 759]. criterion [518]. critical [785]. cube [341, 53]. curvature [528, 721]. curve [166]. curved [308]. curves [728, 87, 433, 478, 578, 703, 684, 511, 753, 20, 372]. cusp [248]. cycles [79, 105, 474]. Cyclic [717, 284, 764]. cyclicly [99]. cyclides [62]. Cyclotomic [206, 130].

d’algèbres [225]. damped [52]. dans [599, 577, 790, 538, 298, 407]. decaying [723, 663]. decomposability [6]. Decomposable [651]. Decomposition [328, 728, 5, 60, 459, 793, 422]. Decompositions [436, 455, 188]. decrease [150]. decreasing [249]. deficient [176]. definability [54]. defining [77]. definite [191]. definition [640, 499, 640]. deformations [139, 402, 402]. degenerate [785]. degré [316]. Degree [293, 105, 41, 490, 27, 316]. demi [301]. demi-groupe [301]. Dempwolff [205]. dendrites [571]. Dense [392, 325, 379, 608, 190]. Densities [782]. d’équations [280, 631]. Derivations [354, 476, 537, 681, 659, 215, 189, 262, 378, 454, 427]. derivatives [589, 752, 574, 23, 721, 522]. derived [229]. descent [645]. Designs [760, 257, 203]. details [193]. determined [177]. d’Euler [44]. deux [198]. Developability [175]. developable [232]. deviations [695]. d’holomorphie [580]. Diagonal [416, 161]. diagonals [326, 579]. diagrams [396]. Diameters [174]. Diamond [147]. difference [388, 655, 443]. differentiable [747]. differential [705, 143, 644, 83, 483, 115, 424, 759, 399]. Differentiation [181, 487, 144]. différentielles [631]. Dilations [188]. dimension [76, 602, 563, 390, 88]. dimensional [181, 41, 193, 500, 497, 673, 167, 687, 649, 431, 532]. diminishing [584]. Dirac [610]. Direct [229, 616]. Directed [529, 12]. directional [23, 522]. Directionally [408]. Dirichlet [295, 209, 505, 565, 44, 225, 519]. discontinuous [115, 443]. Discrete [435, 462, 506, 680, 610]. discreteness [143, 28]. disfocality [483]. disjoint [141, 102]. disk [536]. dissipation [631]. Distance [423, 590]. Distance-genericity [423]. Distinguished [207, 250, 681]. distribution [602, 620, 161]. distributions [540, 757, 376]. distributive [33, 545, 587]. divergence [303]. Divisibility [252, 428]. Divisible [80, 655, 257]. division [278, 366, 206]. divisor [236, 764]. divisors [49, 710]. domain [215, 158]. domaines [538, 298]. domains [675, 252, 30, 545, 489, 95, 250, 211]. dominated [630]. double [777, 541, 359]. Dual [659, 166]. duality [33, 463, 247, 471, 40]. Duffing [631]. d’un [449, 301]. d’une [266]. d’unités [198, 316]. Dupin [62]. dyadic [144]. Dyer [774]. dynamics [264].

edge [274]. effective [314]. eigen [376]. eigen-functions [376]. eigenfunctions [438]. eigenvalue [142]. Eigenvalues [748]. Eilenberg [183]. Eisenstein [248, 319, 391]. Elementary [733, 603, 616]. Elements [306, 215, 192, 227]. elliptic [406, 703, 723, 731, 72, 386, 663, 424]. elliptically [652]. embedded [652]. embedding [75]. Embeddings [168, 671, 53]. empty [643, 718]. End [624, 751]. Endomorphism [413, 172]. endomorphisms [679]. ends [624]. enough [347]. enriched [360]. Ensembles [556]. entire [715, 723, 592, 731, 399]. enumeration [274, 361]. envelope [301]. envelopes [535]. enveloping [502]. enveloppes [580]. Epimorphisms [557]. equal [586]. equality [265]. equation [696, 757, 548, 387, 52]. equations [705, 430, 406, 644, 280, 362, 209, 483, 723, 731, 732, 528, 386, 424, 759, 416, 399, 580]. Equicontinuity [450]. Equivalence [7, 756, 244, 712]. equivalences [684]. equivalent [549]. equivariant [691]. Eratosthenes [660]. Erdmann [42]. Erdos [170]. Ergodic [371, 743, 577, 348, 496, 690, 329, 692, 400, 69]. ergodicity [624]. ergodique [577, 690]. Errata [337]. Erratum [338]. error [517]. Espace [449]. espaces [599, 538, 298]. essential [158, 741]. Essentially [145]. Estimates [104, 528, 368, 424]. Estimating [670]. eta [691]. états [449]. Euclidean [602]. Euler [377, 518, 173, 368, 63, 44]. even [731, 627]. every [3, 61, 596]. Exact [477]. example [193]. exceptional [383]. Excising [594]. Existence [727, 775, 631, 52, 723]. Expanding [782]. expansion [517]. Expansions [152, 312, 332, 186, 453]. experiments [342]. Explicit [641, 654, 169, 459]. exponent [333]. Exponential [264, 715, 567]. exponentials [152]. exponents [369]. expressions [115]. Extending [543, 101]. extensible [99]. extension [736, 327, 217, 494, 39]. extensions [581, 762, 94, 373, 284, 74, 621]. exterior [6, 422]. Extremal [575, 459]. Extreme [500, 13, 385, 58].

Faber [170]. facteur [449]. Factor [243, 449, 3, 61, 120]. Factor-correspondences [243]. Factorization [784, 644, 83, 388]. factors [182, 554, 686, 679, 625, 21]. faibles [225]. faisceaux [470]. Families [782, 264, 450, 304]. family [540, 585, 689]. FC [761, 98]. FC-by-finite [761]. Fejes [435]. Feldheim [494]. fibrations [244]. fibre [244]. field [94, 171, 699, 130]. fields [30, 37, 207, 189, 198, 498, 248, 319, 542, 416, 227]. finies [70]. finitary [463, 484]. Finite [495, 780, 761, 379, 41, 602, 91, 117, 562, 343, 490, 546, 500, 572, 497, 27, 792, 748, 382, 461, 632, 689, 646, 128, 416, 59, 532]. finite-dimensional [41, 500, 497, 532]. finitely [624, 3, 61, 286, 159, 128]. finiteness [277]. five [193]. five-dimensional [193]. Fixed [430, 110, 245, 572, 36, 692, 569]. fixing [746]. FK [635]. FK-spaces [635]. flat [73, 138]. flexible [278]. flow [488]. Fonction [438]. fonctions [538, 698, 225, 438, 744]. fondamentaux [198, 316]. Ford [295]. form [691, 248, 460]. forms [383, 765, 8, 345]. formula [266, 121]. formulas [715, 726, 641, 654]. formule [266]. foundation [508, 467]. founded [365]. Four [595]. Fourier [439, 547, 573]. Fractional [752, 66]. Frame [542]. Fredholm [215, 84]. Free [481, 107, 651, 516, 653, 6, 3, 61, 619, 484, 133, 670, 241, 749, 604, 722, 461, 93, 768, 627]. French [577, 790, 266, 449, 46, 298, 280, 235, 352, 253, 402, 690, 16, 711, 640, 456, 198, 316, 470, 407, 301, 420, 44, 225, 493, 438]. Frobenius [614, 24]. Fuchs [546]. Fuchsian [295, 671, 749]. Fuglede [367]. full [743]. function [783, 320, 140, 720, 248, 319, 249, 438]. functional [38, 101, 460]. functionals [182, 507]. functions [191, 435, 678, 104, 778, 55, 152, 555, 590, 312, 772, 715, 219, 694, 592, 784, 636, 81, 536, 773, 706, 442, 376, 443, 770, 68, 598, 750, 617, 118, 469, 658, 364, 258, 561, 726, 747, 697, 23, 225, 240, 468, 541, 641, 157, 186]. functors [7, 410]. fundamental [387, 702, 198, 316]. Fundamentals [711]. further [336]. fuzzy [600].

Gabriel [413]. Galois [284, 656, 24]. galoisiennes [744]. game [264]. gauche [16]. Gaussian [615, 695]. General [739, 319, 460]. generalisation [126, 266]. Generalization [199, 266, 46, 41, 336, 432, 46]. generalizations [96]. Generalized [678, 209, 34, 530, 501, 791, 368, 23, 767, 522, 453, 281, 757, 215, 377, 334, 387, 173, 411, 462, 506, 658, 364, 573]. generate [484]. generated [3, 61, 151, 128, 661]. generating [186]. generation [593, 626]. Generators [550, 135, 216, 172, 269]. genericity [423]. genus [210, 586]. geodesic [166]. Geometric [473]. geometrically [630, 617]. geometries [527, 102, 167]. geometry [62, 15, 566]. germs [771, 177]. Ginsburg [46]. given [268, 596]. Gleason [140]. Global [386, 727, 171, 108, 248, 319]. Godbillon [665]. Gorenstein [725, 511]. Graded [539, 252, 95, 477]. gradients [522]. grading [1, 124]. graph [112]. graphs [259, 165, 79, 105, 529, 629, 597, 221, 174, 65]. Green [420, 438]. Grothendieck [613]. Group [337, 333, 98, 39, 375, 729, 740, 281, 476, 237, 49, 91, 419, 555, 117, 119, 703, 505, 565, 439, 490, 107, 437, 495, 746, 241, 271, 230, 632, 689, 195, 1, 124, 633, 255, 622, 291]. group-automorphisms [437]. groupe [301]. groupes [70, 16]. Groupoid [360]. Groups [3, 61, 497, 465, 740, 122, 295, 761, 35, 581, 306, 550, 236, 612, 479, 25, 619, 671, 691, 343, 563, 572, 16, 107, 733, 383, 27, 135, 414, 238, 444, 257, 642, 735, 176, 429, 241, 210, 311, 544, 749, 585, 604, 722, 461, 93, 98, 213, 248, 319, 614, 22, 712, 515, 358, 160, 570, 223, 652, 255, 39, 649, 168, 622, 477, 24, 59, 743, 650, 432]. growth [678, 676, 399].

Hahn [188, 793, 64]. Half [296, 618, 565]. half-line [618]. half-space [565]. Half-turns [296]. Hamiltonian [42, 438]. hamiltonien [438]. Hamming [185]. handlebody [48]. Hankel [757, 714, 472]. Hardy [719, 574]. Harish [762].

Harish-Chandra [762]. Harmonic

[675, 565, 555, 736, 253]. harmonics [439]. harmonique [401, 253]. Hausdorff [508, 463, 199, 602, 261]. having [5, 193, 444, 1, 124]. heat [696]. Heisenberg [237, 555, 505, 565, 439, 712]. Helly [58]. Helly-type [58]. Hensel [791]. hereditary [77, 537]. heredity [739]. Hermite [56]. Hermitian [167, 672, 645]. Heyting [434]. High [71, 615, 165]. High-power [71]. Higher [681, 614, 687]. higher-dimensional [687]. highly [221]. Hilbert [748, 364, 249, 667, 325]. Hille [514]. Hjelmslev [429, 201]. Hodge [673]. Hölder [425]. holes [597]. Holland [338]. holomorphe [253, 310]. holomorphes [538, 698]. Holomorphic [240, 783, 778, 253, 706]. homeomorphic [53, 190]. homeomorphism [606]. homeomorphisms [624, 674]. Homogeneous [751, 492, 208, 163]. Homologies [640]. Homology [287, 735, 682]. homomorphic [285]. Homomorphism [370]. Homomorphism-compact [370]. Homomorphisms [587, 750, 429]. homotopie [235]. homotopy [360, 235, 245, 709, 149, 639, 47, 114, 146]. Hopf [350, 601]. hub [235]. hull [281, 509]. hulls [729]. hyperbolic [657, 22]. hypercenter [426]. hypercyclic [491]. hyperelliptic [409]. hypergeometric [777, 658, 541]. Hypermetric [185]. hyperplane [788]. hyperspace [504, 395]. Hyperspaces [85]. hypersurfaces [423, 547].

IA [722]. ideal [30, 593, 626]. idealizers [113]. Ideals [19, 32, 738, 545, 618, 394, 421, 220, 347, 113, 147, 10, 776]. idéaux [599]. Idempotent [666, 301]. Idempotents [634, 49]. Identities [760, 90, 541]. identity [529, 334, 251, 359, 132]. II [738, 182, 696, 67, 623, 551, 273, 101, 656, 737, 215, 505, 169, 371, 322, 765, 586, 499, 338, 319, 26, 179, 718, 307, 743]. III [565, 300, 570]. images [285, 754]. immersions [552]. implicative [267]. implying [90]. Inclusion [635, 57, 299]. Inclusions [693, 42]. Incompressible [48]. indecomposable [294]. independent [498]. index [245, 679]. indexed [12]. indistinguishability [398]. Induced [230, 451, 254, 656, 189]. inducing [160]. Induction [560, 614]. inductive [390, 29]. Inequalities [289, 31, 680, 356, 141, 72]. inequality [425, 71, 305]. infinite [674, 707, 600, 296]. Infinitely [466, 177]. Inflectional [433]. information [387]. infrasequential [407]. infraséquentielles [407]. Injective [122, 2, 123, 294, 509, 646, 100]. Injectives [382]. Injectivity [434]. inner [378, 297]. integers [171, 595, 710]. integrability [475]. integrable [715]. Integral [81, 255, 380, 273, 107, 158, 766, 561]. integrals [554, 765]. integration [34, 358]. Interior [424, 138]. intermediate [207]. interpolation [778, 708, 592, 292, 154, 228]. Intersection [304, 632, 689]. Intersections [70, 37, 546]. intervals [141, 334, 636]. intrinsic [34]. Introduction [758]. Invariance [106, 665]. Invariant [782, 21, 729, 353, 691, 473, 89, 730]. invariants [103, 27]. inverse [31, 777, 301]. inverses [752, 598]. inversive [229]. invertible [354, 347]. inverting [573]. involution [262, 242]. involutorial [481]. involving [485]. Irreducibility [688, 204, 336]. Irreducible [195, 490, 495, 497, 746, 754]. Irregularity [150]. isolated [351]. Isometric [9]. Isometries [131, 211, 302, 690, 469, 258, 690]. Isomorphic [241, 282]. isomorphisms [448]. isotopes [282]. Isotropic [552]. iteration [705]. IV [50, 649, 710].

Jacobi [200, 591, 517, 529, 196, 494, 654]. Jame [86]. James [9]. Jensen [356]. jets [349]. join [465]. Jordan [790, 41, 263, 793].

Kac [323]. Kasparov [477]. kernel [121]. kernels [708]. KK [569]. knot [653]. Knots [133, 17, 466, 756, 355]. Koebe [104, 608]. Korteweg [548]. Köthe [773]. Kowalsky [405]. Krull [414, 236, 311]. Kubert [376]. Kubilius [71].

L. [546]. labeled [202]. labelled [274]. lacunary [693]. Lagrangian [274, 126]. Lambert [57]. Lamperti [371]. Landau [169]. languages [670]. Laplacian [320, 505]. Large [695, 444, 390, 275, 649, 416]. largest [77, 490]. Lashof [774]. Latin [326]. lattice [533, 239, 281, 346, 545, 458, 223]. lattices [92, 73, 33, 286, 277, 55, 448, 14, 159, 484, 587, 558, 417, 692, 128, 120, 787, 231]. Laurent [213]. LCA [176]. left [16, 113, 325]. lemma [38, 791]. length [627]. L’enveloppe [301]. less [586]. leurs [580]. level [615]. Lexicographic [616]. Lie [659, 709, 383, 735, 220, 389, 323, 622, 454]. lifting [331]. lifts [749]. light [742]. like [513]. LIL [460]. Limits [203]. Lindelöf [568]. line [618]. linéaires [631]. Linear [469, 258, 721, 182, 479, 101, 110, 188, 27, 238, 217, 570]. Linearization [700, 200, 196]. linked [256]. links [156, 261]. Lipschitz [522]. Lipschitzian [408]. Local [108, 363, 30, 571]. Local-global [108]. locale [662]. locales [781]. Localisation [456]. Localization [151, 158, 134, 456]. Localizations [256, 95, 351]. Locally [73, 171, 201, 317, 568, 384, 40, 261, 661]. logic [600, 648]. logics [583]. Longest [79, 105]. loop [651, 282]. loops [282]. Lovász [170]. lower [489, 619, 657]. Lucas [297]. Lucas-sets [297]. Lusternik [737]. Lying [164]. Lying-over [164]. Lyndon Macaulay [526, 97]. MacLane [183]. Mahler [638]. major [637]. majorants [308]. manifold [674]. manifolds [624, 606, 751, 492, 657, 542, 63]. many [624, 418, 286]. map [596, 665, 112, 177]. mapping [608, 737]. mappings [675, 408, 232, 441, 36, 462, 506, 389, 163, 256, 569]. maps [87, 202, 188, 183, 351, 361, 197, 686, 293, 192, 125]. Markov [141]. Martin [503]. Massey [464, 619, 47]. matrices [268, 139, 418, 683, 246, 646, 447, 672, 767]. Matrix [630, 268, 113]. matroids [127]. Maximal [709, 394, 172, 113, 740, 562, 311, 632, 689, 323, 130]. maximum [105, 214, 694]. means [473, 81, 573]. measure [380, 674, 707, 531, 565, 329, 283, 638]. measure-preserving [329]. measures [315, 239, 602, 457, 547]. meet [562]. Meixner [266]. Mellin [453]. Mendelsohn [760]. meromorphic [770]. Merotopic [373]. metabelian [39]. metacenter [301]. métacentre [301]. metacompact [384]. metaplectic [391]. method [489, 785]. methods [142, 685, 274, 126]. metric [13, 781, 18, 525, 367]. Metrizability [532]. metrizable [11]. Minimal [115, 464, 417]. minimally [516]. minor [637]. mixed [727, 757, 422]. Mixing [357, 674, 45]. mod [682, 512]. model [543]. Modèles [235]. models [235]. modified [272]. Modular [446, 8, 59, 740, 659, 787]. Modularity [458, 90]. module [6]. modules [451, 762, 294, 669, 195, 687, 100, 622]. modulo [485]. moment [508, 199]. monodromy [210]. monomial [511]. Monomorphisms [571]. Monotonic [782]. Moody [323]. Moore [147]. Morita [712]. most [139]. motion [460]. mots [556]. Moyennes [744]. MR [338]. multiadjoint [410]. multidimensional [496]. multiparameter [487, 692]. multiple [664, 486]. Multiplication [755]. Multiplicative [636, 642]. multiplicité [631]. multiplicity [526, 743]. multiplier [497]. Multipliers [611, 713, 184, 666]. multivalent [617]. mutants [756].

Naturally [492]. nearest [775]. Necessary [143, 265, 228]. negative [191]. Nemyckii [599]. nest [669]. net [193]. nets [70]. Neumann [696, 458, 609, 507, 488, 427]. neutral [732]. Newman [506, 462]. Newton [791]. Nilpotent [581, 515, 160, 323, 622, 318]. no [111]. nodal [344, 438]. nodales [438]. nodes [753]. Noetherian [151, 421, 276, 180, 100, 403]. non [421, 631]. non-linéaires [631]. non-Noetherian [421]. Nonamphicheiral [17]. Noncommutative [322, 41, 784, 224, 717, 687]. nonconnected [612]. nonconvex [23]. nondegenerate [261]. nonessential [158]. nonexistence [792]. nonexpansive [36, 569]. Nonextendability [68]. nonhomogeneous [783]. nonisomorphic [361]. nonlinear [406, 142, 489, 746, 52]. nonlocally [535]. nonmeasurable [54]. Nonnegative [406, 688, 121, 196]. nonoscillatory [362]. nonreflexive [111]. nonresidues [518]. nonsmooth [700]. Nonstandard [618, 40]. Nonzero [76]. norm [78]. Normal [384, 411, 745, 409, 441, 147, 317, 734, 569, 449]. normality [367]. normalizing [151]. normaux [449]. normed [110, 469]. norms [353, 442, 672]. North [338]. North-Holland [338]. note [330, 95, 428, 526, 544, 512, 225, 564, 129]. Notions [711]. nouveaux [470]. nowhere [747]. number [306, 699, 71, 117, 368, 498, 161, 130, 125]. numbers [377, 332, 334, 173, 279, 309]. Nussbaum [245].

objects [352, 316]. obstacles [247, 471]. obtained [540]. Obtaining [400]. occupation [615]. odd [167, 88]. odd-dimensional [167]. one [142, 35, 751, 293, 21, 431]. one-dimensional [431]. one-relator [35]. ones [484]. onto [675, 557]. Open [197, 462, 506]. Opérateur [298, 599]. opérateurs [577]. operational [583]. Opérations [493, 287, 493]. operator [298, 353, 220, 378, 313, 129, 743, 427]. operator-valued [427]. Operators [607, 120, 143, 603, 577, 270, 83, 388, 237, 115, 145, 741, 452, 450, 371, 574, 748, 588, 84, 755, 514, 488, 734]. optimal [42]. Optimization [483]. orbit [479, 444]. orbit-stabilizer [479]. Orbital [765]. orbits [540, 329, 544]. Order [55, 533, 67, 344, 306, 440, 731, 732, 792, 45, 759, 431]. Order-Cauchy [55]. ordered [281, 358, 217, 223]. orderings [745, 50, 363]. orders [365, 311]. Ordinary [249, 83, 483]. Orlicz [590, 615, 695, 569]. ortho [234]. ortho-bases [234]. orthocomplemented [787]. orthocompletion [281]. Orthogonal [440, 551, 135, 89, 271, 748, 486, 561, 203]. Orthogonality [219]. orthogroups [194]. Orthomodular [286, 277]. Oscillation [72, 759, 84]. Oscillations [732]. Outer [454, 108]. Ovoids [321]. Oxford [338].

packings [346]. pair [777]. Pairs [32, 738, 164, 663, 59]. pairwise [203]. paracompact [384]. Paracompactness [568]. parameter [142]. paramètres [662]. Parseval [757]. part [81, 240]. partial [644, 424]. partially [358]. parties [420]. partition [274, 160]. partition-inducing [160]. Partitioning [527]. partitions [251]. parts [251, 140, 420]. paths [615, 695]. patterns [304]. Perfect [291, 754, 192]. period [133]. periodic [466, 507, 488]. periodicity [582]. périodiques [631]. permutation [343, 544]. perturbations [178, 21]. Pettis [475]. phase [785]. Picard [119]. Piecewise [782]. Pixley [384]. Planar [202, 361, 125]. Plane [753, 788, 578]. planes [495, 444, 205, 321, 792, 201, 88, 229, 649]. plans [538]. plaque [235]. Plessner [728]. plurisubharmonic [694]. plusieurs [662]. poids [298]. Poincaré [166, 97]. point [110, 245]. points [430, 781, 775, 58, 785, 335, 703, 18, 36, 692, 593, 626, 569]. Pointwise [701, 329, 443]. Poisson [380, 121]. Polar [455, 193, 482]. polycyclic [195]. polyhedra [510, 193]. polyhedrons [340]. Polynomial [729, 27, 213, 275]. Polynomials [366, 56, 289, 369, 288, 551, 440, 688, 517, 716, 292, 748, 308, 179, 300, 459, 575, 121, 200, 196, 494, 654, 297]. ponctuel [577]. Porosity [589]. poset [509, 754]. positive [727, 315, 708, 723, 500, 81, 271, 386, 663, 712, 459, 240]. Postnikov [628]. potentials [320]. pour [599, 662, 711]. Power [176, 577, 71, 539]. power-bounded [577]. power-deficient [176]. Power-rich [176]. powers [272, 150, 226]. pramarts [153, 480]. pre [429]. pre-Hjelmslev [429]. preduals [9]. preference [365]. prescribed [289, 105, 326, 400]. presented [159]. preserve [303]. preserving [674, 707, 87, 329, 686]. presheaves [639]. presque [44]. Prime [334, 32, 738, 262, 421, 576, 521, 355, 485]. primes [151, 226]. Primitive [269, 74]. principal [30, 115]. principles [110, 694, 147]. Pro [410]. Pro-categories [410]. Problem [337, 696, 199, 479, 505, 565, 194, 546, 180, 472, 179, 300, 275, 787, 147, 20]. problems [727, 647, 142, 489, 555, 338, 663, 459, 519]. processes [181, 487, 496, 615, 695]. processus [662]. product [774, 339, 77, 108, 352, 583, 229, 121, 200, 196, 726, 392, 10, 297]. Products [419, 271, 734, 761, 464, 619, 563, 397, 702, 47, 44]. Produit [352]. produits [44]. projectable [281]. projection [78]. Projections [513, 538, 458, 661, 776]. Projective [102, 687, 412, 192, 330, 788, 51, 527, 495, 601, 238, 167, 792, 327, 649]. projectives [382]. projectivity [187]. proof [38, 435]. proper [197]. Properties [268, 455, 56, 533, 252, 603, 385, 137, 18, 80, 473, 276, 558, 371, 640, 764, 173, 327, 575, 44, 514, 569]. Property [315, 5, 245, 331, 443, 462, 506, 367, 395, 465]. propres [438]. Propriétés [44, 640]. Proximal [553, 623]. proximities [260]. proximity [390]. pseudocharacter [75]. Pseudocomplemented [267]. Puiseux [791]. puissances [577]. Pure [30, 546, 526].

quadratic [198, 345, 621]. quadratiques [198]. quadrics [193]. quantificational [704]. quasi [470, 44]. quasi-sûres [44]. quasi-sure [44]. quasi-unipotent [470]. quasi-unipotents [470]. quasicentral [359, 132]. quasifields [642]. Quasigroup [760]. quasireflexive [86]. quasisemilocal [216]. quasisimilar [741]. quasisimilarity [647]. quasizeroes [43]. quaternary [485]. quaternion [230]. quaternionic [256]. quintic [518]. quotient [91, 172, 166, 180]. quotients [595].

Racah [121]. Radial [731]. radical [618, 739]. radii [340]. Radius [407]. radix [332]. Radon [742]. raising [129]. Ramanujan [96]. Random [551, 174]. Range [442, 457]. rank [786, 709, 779, 525, 461, 323, 21]. rank-metric [525]. ranking [157]. Rankings [157]. rate [150]. rates [303, 400]. Rational [244, 26, 31, 312, 235, 703, 561]. rationnelle [235]. Rayon [407]. rays [13]. Real [278, 779, 423, 37, 772, 81, 140, 720, 22, 240, 415, 130]. realization [293, 355]. really [224]. Rearrangements [303]. reciprocity [309]. reconnaissables [556]. Recursive [221, 65]. reducibility [541]. reducible [99, 338, 313]. reductive [612, 492, 248, 319]. reflection [25]. reflections [419, 271]. Reflexive [769, 178, 629, 597]. reflexivity [222]. regions [295, 298, 4]. Regular [771, 239, 243, 79, 563, 226, 763, 163]. regularité [407, 698]. regularity [175, 445, 405, 407]. related [705, 714, 443]. relating [251]. relation [57]. Relations [638, 420]. relative [655, 456, 456]. Relativized [45]. relator [35]. remarks [722, 391]. remarque [225]. remote [335, 18]. repetitions [251]. Representation [715, 155, 389, 633, 196, 561]. Representations [467, 365, 605, 254, 613, 762, 601, 497, 702, 59, 485]. residuacity [485]. residue [656]. resolution [396, 280]. resolutions [526]. résolvantes [662]. Resolvent [573, 209]. respect [219, 274]. respecting [653]. restricted [31, 486]. restriction [560]. restrictions [587]. result [126, 627]. Résultats [470]. results [272, 237, 470, 398]. retracts [629, 148]. revisited [287, 118]. Rewritable [761]. Riccati [72]. Rich [260, 176]. Riemann [724, 401, 580, 34, 210, 218]. Riemannian [492]. Riesz [394]. right [545, 151, 403]. rigid [227]. ring [375, 366, 171, 699, 426, 216]. Rings [294, 347, 32, 738, 243, 545, 328, 49, 333, 95, 55, 164, 108, 656, 725, 563, 109, 107, 151, 262, 421, 135, 172, 491, 414, 276, 413, 241, 534, 285, 201, 324, 98, 242, 389, 213, 180, 74, 646, 255, 539, 39, 97, 403, 113, 227]. Rogers [716]. Root [246, 668]. rooted [637]. roots [418, 154]. Rose [46]. Rotundity [773]. Roy [384]. Ruled [586]. rules [350]. Runge [211].

S. [305]. sûres [44]. same [32, 738]. sample [615, 695]. satisfiability [275]. saturation [363, 10]. Scalar [381]. scattering [548]. schematic [233]. scheme [311]. schemes [543]. Schnirel’mann [737]. Schrödinger [115]. Schrödinger-type [115]. Schur [230]. Schwartz [758]. Secant [372]. second [104, 732, 759, 29]. sectional [586]. sections [788, 409]. segment [5]. selection [38, 504]. self [193, 646]. self-injective [646]. self-polar [193]. Selmer [703]. semi [774, 215]. semi-Fredholm [215]. semi-tensor [774]. Semicontinuity [713]. Semidirect [339]. semigroup [77, 109, 301, 570, 420]. Semigroupe [420]. Semigroups [43, 533, 508, 467, 236, 481, 571, 763, 679]. semihereditary [216]. semilattices [267, 446, 521]. Semilinear [663, 731, 72, 386, 519]. semimartingales [212]. semimodules [411]. semiplanes [655]. Semiprime [285]. Semisimple [600]. semitopological [7]. separable [335, 650]. Separating [232]. sequence [620, 524, 569]. sequences [701, 150, 357, 151, 368, 69, 477, 503, 186]. Serial [403, 285]. series [152, 619, 630, 777, 486, 248, 319, 391, 44, 638, 539, 97, 541, 44]. set [92, 600, 18, 343, 572, 197]. set-stabilizers [343]. set-valued [197]. Sets [434, 729, 553, 623, 232, 5, 58, 51, 693, 28, 54, 562, 473, 655, 304, 442, 12, 217, 404, 186, 291, 297]. several [142, 441]. Shape [684]. sharply [14]. sheaves [122, 2, 123]. Shifts [625]. Shih [640]. short [334, 636]. shuffles [768]. sided [83]. sieve [660]. sieved [551]. Signatures [156, 780]. Simple [261, 646, 87, 41, 481, 383, 16]. simples [16]. Simplicial [345, 103, 639, 114]. simultaneous [549, 161]. singular [380, 143, 186]. singularities [578, 249, 396]. Singularity [511]. six [435, 586]. size [214, 670]. skew [51]. slope [87, 444]. slope-preserving [87]. slow [678, 84]. Small [559, 91, 779, 564]. smallest [76, 576]. Smooth [415, 547, 112]. smoothness [239]. Sobolev [298, 590]. socle [790]. soluble [652]. Solution [546]. solutions [430, 406, 727, 489, 723, 731, 528, 631, 386, 663, 212, 161, 399, 290, 307, 519]. Solvable [318, 444, 432]. Solving [280]. Some [705, 56, 533, 605, 247, 471, 277, 455, 237, 335, 137, 18, 377, 173, 722, 391, 447, 96, 398, 104, 489, 175, 714, 443, 316, 213, 469, 494, 726, 255, 100, 654]. sous [70]. sous-groupes [70]. space [86, 628, 449, 433, 478, 553, 623, 51, 758, 602, 77, 659, 110, 18, 565, 691, 394, 596, 748, 754, 667, 147]. Spaces [183, 596, 50, 393, 651, 513, 315, 185, 270, 370, 568, 463, 635, 408, 775, 232, 9, 379, 78, 111, 373, 384, 298, 75, 335, 409, 590, 153, 209, 85, 116, 40, 457, 500, 784, 501, 482, 261, 535, 773, 390, 224, 322, 719, 211, 302, 615, 695, 574, 148, 469, 424, 258, 472, 780, 661, 524, 363, 564, 392, 317, 297, 569, 532, 372]. spacetimes [673]. Special [320, 729, 437, 560, 447, 650]. specialised [193]. specified [673]. spectra [741, 610, 89, 588]. Spectral [455, 299, 742, 476, 215, 530, 313]. spectrum [143, 283]. sphere [138, 643, 718]. spheres [346, 47, 146]. Spherical [439, 735, 162, 691]. spine [341]. spiral [598]. Spline [118, 584]. splines [523, 118]. split [59]. spreads [238, 20]. squares [326]. Stability [231, 363]. stabilizer [479]. stabilizers [343]. Stable [639, 402, 47]. stables [402]. Starlike [772, 658, 468]. starshaped [58]. state [449]. states [594]. stationary [785, 648]. Steenrod [774]. Steiner [214, 229, 233]. step [29]. Stickelberger [284]. stochastic [615, 695]. Stokes [298]. Stone [80]. strength [739]. stretched [97]. strictly [315]. Strong [153, 712, 503, 182, 424, 524]. strongly [281, 495, 52]. Structural [603, 558, 18]. Structure [100, 159, 397, 669, 776, 569]. study [473]. subalgebras [537, 609, 325]. subbase [64]. Subclasses [468]. subelliptic [505]. subfield [130]. subfields [207, 681]. subgradients [23]. subgraph [275]. subgroup [3, 61]. subgroups [119, 709, 680, 546, 107, 544, 89, 632, 689, 22, 652, 318]. subharmonic [694]. subharmonicity [283]. sublinear [727, 723]. sublinear-superlinear [727]. submanifolds [415]. submeasures [436]. subnormal [588, 465]. subordinate [468]. subpramarts [480]. subsemilattice [223]. subsequential [348]. subset [608]. subsets [379]. subspaces [51, 448, 719, 730, 677, 392, 190, 11]. Sufficiency [591, 349]. sufficient [143, 265]. sum [595, 494, 96]. summability [182, 554, 685, 630, 57, 486]. sums [616, 710]. superadditive [487, 496]. supercuspidal [605]. superlinear [727, 759]. supersolvable [515]. Supplements [425]. sure [44, 44]. surface [409]. surfaces [724, 67, 344, 401, 208, 210, 586, 48, 218, 438]. surgery [582]. switching [606]. Sylow [689]. Sylvester [20]. Symbole [310]. Symbolic [226]. symmetric [426, 766, 524]. symmetry [76, 440]. synthesis [313]. system [214, 280, 498]. système [280]. Systèmes [198, 316]. systems [668, 77, 700, 610, 198, 316, 246, 229, 233].

Tamarkin [514]. Tangent [409, 138, 704]. tangential [476]. tangles [355]. Tauberian [685, 567]. Tchebycheff [280]. Tchebychev [575]. technique [484]. Tensor [563, 702, 774, 352, 397, 583]. tensoriel [352]. tensors [721]. Term [144, 382]. test [647]. th [154]. their [467, 388, 288, 437, 414, 749, 748, 712, 721, 453]. theorem [577, 508, 478, 685, 58, 204, 379, 46, 496, 504, 737, 336, 567, 690, 262, 169, 66, 329, 584, 400, 69, 726, 314, 293, 622, 645, 432, 64]. théorème [577, 46, 690]. Théorèmes [599, 662]. theorems [348, 635, 31, 692, 399]. théorie [456, 493]. theories [451]. Theory [553, 705, 600, 31, 60, 71, 656, 360, 215, 40, 530, 673, 149, 639, 791, 456, 683, 411, 648, 665, 633, 789, 679, 621, 667, 645, 493, 64]. theta [676]. Thrall [29]. Three [647, 276, 399, 67, 344, 440, 332]. tightness [239]. tiling [670]. tilting [451]. timely [577]. times [615]. Toeplitz [270, 84]. Topological [25, 764, 327, 508, 247, 471, 473, 276, 582, 352].

Topologically [139, 7].

topologically-algebraic [7]. topologie [253]. topologies [171, 413]. topologique [352]. topology [253, 34, 524]. Topos [434]. toposes [613]. tori [717, 687]. Torsion [451, 106, 703, 456, 461]. torsion-free [461]. torsionfree [172]. total [187]. totally [708, 415]. Tóth [435]. tournante [235]. trace [101, 549]. Traces [263]. trajectories [700]. transcendental [399]. transferable [14]. transform [439, 364, 742]. transformable [472]. transformation [757, 548, 329]. Transformations [782, 707, 714, 329, 72, 82]. transforms [708, 547, 573]. transitive [259]. Transitivity [234]. Translation [88, 444, 205, 321, 342]. transversal [214]. transverse [728]. tree [513]. tree-like [513]. trees [202, 274, 126, 637]. triangularizability [549]. Tricomi [266]. trinomials [4]. triple [214]. Trivial [343, 19]. Trudi [529]. Tubes [676]. Turán [71, 179, 300]. Turán-Kubilius [71]. turns [296]. Tutte [126]. twelve [595]. twisted [213, 621]. twists [703]. Two [385, 272, 139, 83, 41, 27, 673, 777, 198, 88]. two-dimensional [673]. two-sided [83]. type [449, 685, 58, 94, 115, 715, 567, 136, 383, 646, 424, 679, 625, 427, 396].

ultraspherical [288]. Unbounded [191, 489, 452, 588]. uncountable [45]. undecidability [272]. unicité [631]. unified [445]. Uniform [369, 736, 112, 138, 517, 283, 69]. uniformity [393]. Uniformly [661]. unions [102, 11]. unipotent [531, 470]. unipotents [470]. unique [393]. Uniqueness [162, 653, 47, 375, 673, 186, 291]. uniserial [294]. unit [563, 107, 98]. unit-regular [563]. unital [30]. Unitarily [353]. unitary [729, 633]. units [333, 198, 316, 498]. unity [154]. Univalent [658, 104, 772, 4]. Universal [475, 228]. universe [317]. unsaturated [59]. unsymmetric [209]. upper [489, 742]. ups [788]. use [597]. using [272, 229].

valeurs [744]. Valuation [227, 135]. valuations [745]. value [489, 555, 576]. valued [600, 480, 352, 219, 434, 773, 197, 460, 427, 297]. values [354, 784, 638, 415, 352]. vanishing [528]. variable [369, 777]. variables [441]. variants [445]. variation [678, 608, 584]. variation-diminishing [584]. Variational [142]. variations [42, 591, 503]. varieties [533, 629, 417, 382, 223]. variety [531]. Vaughan [334]. vector [73, 379, 55, 480, 457, 690, 773, 297]. vector-valued [773, 297]. vectoriel [690]. versal [139, 402]. verselles [402]. version [314]. vertex [259, 478]. vertex-transitive [259]. vertical [431]. Vey [665]. view [445]. VMOA [724]. Volterra [150, 452]. volume [657]. Voronovskaya [584]. Vries [548].

walk [551]. Wallman [155]. wave [528, 52]. Weak [674, 212, 476, 457, 45, 225]. Weakly [770, 427, 406, 74]. wedges [698]. Weierstrass [349]. Weighted [348, 719, 78, 302]. weights [298, 219]. well [365, 202]. well-founded [365]. Weyl [439]. Which [217, 3, 61, 286, 111, 282, 473, 596, 444, 120]. Whitehead [614]. Whitney [684]. whose [294, 329, 595, 285]. width [748, 128]. Wilson [716, 121]. Witt [328, 699, 725]. Word [338, 194, 787, 627]. word-length [627]. words [768].

Young [129].

Zassenhaus [454]. Zero [49, 289, 595].

References

Orzech:1976:BGA

[1] Morris Orzech. On the Brauer group of algebras having a grading and an action. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):533–552, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [124].

Dobbs:1977:CIS

[2] David E. Dobbs. On characterizing injective sheaves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1031–1039, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [123].

Brunner:1979:GWE

[3] A. M. Brunner and R. G. Burns. Groups in which every finitely generated subgroup is almost a free factor. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 1329–1338, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigenda [61].

Rahman:1980:CRU

[4] Q. I. Rahman and J. Waniurski. Coefficient regions for univalent trinomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1–20, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Breen:1980:DSH

[5] Marilyn Breen. A decomposition for sets having a segment convexity property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):21–26, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boratynski:1980:CDE

[6] M. Boratyński, E. D. Davis, and A. V. Geramita. Complete decomposability in the exterior algebra of a free module. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):27–33, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herrlich:1980:ETA

[7] H. Herrlich, R. Nakagawa, G. E. Strecker, and T. Titcomb. Equivalence of topologically-algebraic and semitopological functors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):34–39, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maher:1980:MFC

[8] David P. Maher. Modular forms from codes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):40–58, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1980:IPJ

[9] Leon Brown and Takashi Itô. Isometric preduals of James spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 59–69, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wagon:1980:SPI

[10] Stanley Wagon. The saturation of a product of ideals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):70–75, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanDouwen:1980:UMS

[11] E. K. van Douwen, D. J. Lutzer, J. Pelant, and G. M. Reed. On unions of metrizable subspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):76–85, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Millet:1980:CCA

[12] Annie Millet and Louis Sucheston. Convergence of classes of amarts indexed by directed sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):86–125, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Avis:1980:ERM

[13] David Avis. On the extreme rays of the metric cone. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(1):126–144, February 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gratzer:1980:CST

[14] G. Grätzer and C. R. Platt. A characterization of sharply transferable lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):145–154, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Perkel:1980:CG

[15] Manley Perkel. A characterization of PSL(2, 31) and its geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 155–164, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Handelman:1980:ASG

[16] David Handelman. Algèbres simples de groupes à gauche. (French) [Simple algebras of left groups]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):165–184, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Acuna:1980:NCK

[17] F. González Acuña and José M. Montesinos. Nonamphicheiral codimension 2 knots. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):185–194, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gates:1980:SSP

[18] Catherine L. Gates. Some structural properties of the set of remote points of a metric space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):195–209, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Geramita:1980:ITC

[19] A. V. Geramita and C. A. Weibel. Ideals with trivial conormal bundle. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):

210–218, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Watson:1980:SPS

[20] Kym S. Watson. Sylvester’s problem for spreads of curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):219–239, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thompson:1980:IFU

[21] Robert C. Thompson. Invariant factors under rank one perturbations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 240–245, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

OMalley:1980:SRH

[22] Thomas J. O’Malley. S-subgroups of the real hyperbolic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):246–256, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rockafellar:1980:GDD

[23] R. T. Rockafellar. Generalized directional derivatives and subgradients of nonconvex functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):257–280, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sonn:1980:FGG

[24] Jack Sonn. SL(2, 5) and Frobenius Galois groups over Q. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):281–293, ????

1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Djokovic:1980:TRG

[25] Dragomir Ž. – Doković. Topological reflection groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):294–309, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Newman:1980:RAI

[26] D. J. Newman and A. R. Reddy. Rational approximation to x^{n}. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 310–316, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Huffman:1980:PIF

[27] W. Cary Huffman. Polynomial invariants of finite linear groups of degree two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):317–330, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Freiwald:1980:DAS

[28] Ronald C. Freiwald. k-discreteness and k-analytic sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):331–341, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smalo:1980:ISS

[29] Sverre O. Smalø. The inductive step of the second Brauer–Thrall conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):342–349, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benabdallah:1980:PUL

[30] K. Benabdallah and K. W. Roggenkamp. Pure unital local principal ideal domains in local fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):350–353, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1980:IIT

[31] Peter Borwein. Inequalities and inverse theorems in restricted rational approximation theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):354–361, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1980:PRS

[32] David F. Anderson and David E. Dobbs. Pairs of rings with the same prime ideals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):362–384, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Banaschewski:1980:DDC

[33] B. Banaschewski. The duality of distributive continuous lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 385–394, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Henstock:1980:GRI

[34] Ralph Henstock. Generalized Riemann integration and an intrinsic topology.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):395–413, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brunner:1980:COR

[35] A. M. Brunner. On a class of onerelator groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):414–420, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lim:1980:ACF

[36] Teck Cheong Lim. On asymptotic centers and fixed points of nonexpansive mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):421–430, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Craven:1980:IRC

[37] Thomas C. Craven. Intersections of real closed fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):431–440, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baggett:1980:FAP

[38] L. W. Baggett and Arlan Ramsay. A functional analytic proof of a selection lemma. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):441–448, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roggenkamp:1980:GRM

[39] K. W. Roggenkamp. Group rings of metabelian groups and extension categories. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):449–459, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grainger:1980:NDT

[40] Arthur D. Grainger. On the nonstandard duality theory of locally convex spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):460–479, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Conlon:1980:GDT

[41] Mary Ellen Conlon. A generalization of degree two simple finitedimensional noncommutative Jordan algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):480–493, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clarke:1980:ECH

[42] Frank H. Clarke. The Erdmann condition and Hamiltonian inclusions in optimal control and the calculus of variations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):494–510, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rankin:1980:SQ

[43] S. A. Rankin and C. M. Reis. Semigroups with quasizeroes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):511–530, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Queffelec:1980:PPS

[44] Hervé Queffélec. Propriétés presque sûres et quasi-sûres des séries de Dirichlet et des produits d’Euler. (French) [Almost-sure and quasi-sure properties of Dirichlet series and Euler products]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):531–558, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McMahon:1980:RWM

[45] Douglas McMahon. Relativized weak mixing of uncountable order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 559–566, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chaput:1980:GDT

[46] Daniel Chaput and Gert Sabidussi. Une généralization du théorème de Ginsburg–Rose. (French) [A generalization of the Ginsburg–Rose theorem]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):567–575, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kochman:1980:UMP

[47] Stanley O. Kochman. Uniqueness of Massey products on the stable homotopy of spheres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):576–589, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lyon:1980:ISB

[48] Herbert C. Lyon. Incompressible surfaces in the boundary of a handlebody— an algorithm. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):590–595, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cliff:1980:ZDI

[49] Gerald H. Cliff. Zero divisors and idempotents in group rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):596–602, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marshall:1980:SOI

[50] Murray Marshall. Spaces of orderings. IV. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):603–627, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruen:1980:BSS

[51] Aiden A. Bruen. Blocking sets and skew subspaces of projective space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 628–630, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Webb:1980:EAB

[52] G. F. Webb. Existence and asymptotic behavior for a strongly damped nonlinear wave equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):631–643, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartman:1980:HEC

[53] Jehuda Hartman. On homeomorphic embeddings of K_{m,n} in the cube. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 644–652, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Friedman:1980:DNS

[54] Harvey Friedman. On definability of nonmeasurable sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):653–656, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dashiell:1980:OCC

[55] F. Dashiell, A. Hager, and M. Henriksen. Order-Cauchy completions of rings and vector lattices of continuous functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):657–685, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Allaway:1980:SPH

[56] Wm. R. Allaway. Some properties of the q-hermite polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):686–694, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gawronski:1980:IRA

[57] W. Gawronski, H. Siebert, and R. Trautner. An inclusion relation for abel, borel, and Lambert summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):

695–702, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Breen:1980:EPH

[58] Marilyn Breen. (d − 2)-extreme points and a helly-type theorem for starshaped sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):703–713, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tinberg:1980:MRF

[59] N. B. Tinberg. Modular representations of finite groups with unsaturated split (B, N)-pairs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):714–733, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cunningham:1980:CDT

[60] William H. Cunningham and Jack Edmonds. A combinatorial decomposition theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):734–765, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brunner:1980:CGW

[61] A. M. Brunner and R. G. Burns. Corrigenda: “Groups in which every finitely generated subgroup is almost a free factor”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):766, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [3].

Cecil:1980:CGC

[62] Thomas E. Cecil and Patrick J. Ryan. Conformal geometry and the cyclides of Dupin. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):767–782, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Penna:1980:ECC

[63] Michael A. Penna. Euler classes of combinatorial manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):783–803, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vandeVel:1980:HBT

[64] M. van de Vel. A Hahn–Banach theorem in subbase convexity theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 804–820, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schmerl:1980:RCG

[65] James H. Schmerl. Recursive colorings of graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):821–830, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Katchalski:1980:BFT

[66] Meir Katchalski. Boxes in R^{n}—a “fractional” theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):831–838, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bisztriczky:1980:BSO

[67] Tibor Bisztriczky. Biplanar surfaces of order three. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):839–866, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levy:1980:NBC

[68] Ronnie Levy. Nonextendability of bounded continuous functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 867–879, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Olsen:1980:AET

[69] James H. Olsen. Akcoglu’s ergodic theorem for uniform sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 880–884, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benabdallah:1980:IFS

[70] Khalid Benabdallah and Serge Robert. Intersections finies de sous-groupes nets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):885–892, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elliott:1980:HPA

[71] P. D. T. A. Elliott. High-power analogues of the Turán-Kubilius inequality, and an application to number theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):893–907, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Noussair:1980:OSE

[72] E. S. Noussair and C. A. Swanson. Oscillation of semilinear elliptic inequalities by Riccati transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 908–923, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1980:LFV

[73] Marlow Anderson. Locally flat vector lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):924–936, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nicholson:1980:EWP

[74] W. K. Nicholson, J. F. Watters, and J. M. Zelmanowitz. On extensions of weakly primitive rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):937–944, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dow:1980:AES

[75] Alan Dow. Absolute C-embedding of spaces with countable character and pseudocharacter conditions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 945–956, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chan:1980:NSC

[76] G. H. Chan and M. H. Lim. Nonzero symmetry classes of smallest dimension. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):957–968, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Delfour:1980:LCH

[77] M. C. Delfour. The largest class of hereditary systems defining a C_{0} semigroup on the product space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 969–978, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burbea:1980:BPW

[78] Jacob Burbea. The Bergman projection on weighted norm spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 979–986, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bondy:1980:LCCa

[79] J. A. Bondy and M. Simonovits. Longest cycles in 3-connected 3-regular graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):987–992, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Glasner:1980:DPS

[80] S. Glasner. Divisible properties and the Stone–Čech compactification. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):

993–1007, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Holland:1980:IMF

[81] F. Holland and J. B. Twomey. Integral means of functions with positive real part. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1008–1020, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1980:T

[82] P. G. Rooney. On the Y_{u} and H_{u} transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1021–1044, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Browne:1980:TSF

[83] Patrick J. Browne and Rodney Nillsen. The two-sided factorization of ordinary differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1045–1057, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Power:1980:FTO

[84] S. C. Power. Fredholm Toeplitz operators and slow oscillation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 1058–1071, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Friedler:1980:HCS

[85] L. M. Friedler, R. F. Dickman, Jr., and R. L. Krystock. Hyperspaces of H-closed spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1072–1079, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Andrew:1980:JQB

[86] Alfred D. Andrew and William L. Green. On james’ quasireflexive Banach space as a Banach algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1080–1101, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bisztriczky:1980:CSP

[87] Tibor Bisztriczky and Ivan Rival. Continuous, slope-preserving maps of simple closed curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1102–1113, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ostrom:1980:TPD

[88] T. G. Ostrom. Translation planes of dimension two with odd characteristic. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):1114–1125, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lindner:1980:SCI

[89] C. C. Lindner, R. C. Mullin, and D. G. Hoffman. The spectra for the conjugate invariant subgroups of n^{2} × 4 orthogonal arrays. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1126–1139, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Day:1980:CII

[90] Alan Day and Ralph Freese. A characterization of identities implying congruence modularity. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1140–1167, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davitt:1980:AGF

[91] Richard M. Davitt. On the automorphism group of a finite p-group with a small central quotient. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 1168–1176, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Adams:1980:CSL

[92] M. E. Adams and J. Sichler. Cover set lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1177–1205, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mekler:1980:HCA

[93] Alan H. Mekler. How to construct almost free groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1206–1228, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chang:1980:CCF

[94] S. M. Chang and R. Foote. Capitulation in class field extensions of type (p, p). Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1229–1243, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Daepp:1980:NBR

[95] U. Daepp and A. Evans. A note on Buchsbaum rings and localizations of graded domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1244–1249, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ramanathan:1980:SGR

[96] K. G. Ramanathan and M. V. Subbarao. Some generalizations of Ramanujan’s sum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1250–1260, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sally:1980:PSS

[97] Judith D. Sally. The Poincaré series of stretched Cohen–Macaulay rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 1261–1265, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Merklen:1980:GRF

[98] H. Merklen and César Polcino Milies. Group rings over Z_{(p)} with FC unit groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1266–1269, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehman:1980:CCE

[99] B. Lehman. K-coherence is cyclicly extensible and reducible. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1270– 1276, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sarath:1980:SSN

[100] B. Sarath. Structure of some Noetherian injective modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1277–1287, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elliott:1980:ETL

[101] George A. Elliott. On extending the trace as a linear functional. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 1288–1298, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kestenband:1980:PGD

[102] Barbu C. Kestenband. Projective geometries that are disjoint unions of caps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):1299–1305, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1980:AIS

[103] M. Brown and A. G. Wasserman. Arithmetic invariants of simplicial complexes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):1306–1310, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bshouty:1980:EKC

[104] D. Bshouty, W. Hengartner, and G. Schober. Estimates for the Koebe constant and the second coefficient for some classes of univalent functions.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):1311–1324, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bondy:1980:LCCb

[105] J. A. Bondy and R. C. Entringer. Longest cycles in 2-connected graphs with prescribed maximum degree. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 1325–1332, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chapman:1980:ITB

[106] T. A. Chapman. Invariance of torsion and the Borsuk conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 1333–1341, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartley:1980:FSU

[107] B. Hartley and P. F. Pickel. Free subgroups in the unit groups of integral group rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1342–1352, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Estes:1980:LGC

[108] Dennis R. Estes and Jacob R. Matijevic. Local-global criteria for outer product rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1353–1360, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hardy:1980:ASR

[109] Bonnie R. Hardy and Thomas S. Shores.

Arithmetical semigroup rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 1361–1371, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fournier:1980:FPP

[110] Gilles Fournier. Fixed point principles for cones of a linear normed space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 1372–1381, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Casazza:1980:NBS

[111] P. G. Casazza, Bor Luh Lin, and R. H.

Lohman. On nonreflexive Banach spaces which contain no c_{0} or l_{p}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 1382–1389, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weinstock:1980:UAG

[112] Barnet M. Weinstock. Uniform approximation on the graph of a smooth map in C^{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1390–1396, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stone:1980:MLI

[113] David R. Stone. Maximal left ideals and idealizers in matrix rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):

1397–1410, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lakser:1980:HSA

[114] Harry Lakser. The homotopy of simplicial algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1411–1422, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Faierman:1980:MOS

[115] M. Faierman and I. Knowles. Minimal operators for Schrödinger-type differential expressions with discontinuous principal coefficients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1423–1437, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Garcia-Maynez:1980:CCB

[116] A. García-Máynez. Concerning σconnectedness of Baire spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 1438–1447, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Exarchakos:1980:NAF

[117] Theodoros Exarchakos. On the number of automorphisms of a finite p-group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):1448–1458, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Micchelli:1980:SFC

[118] Charles A. Micchelli and A. Sharma. Spline functions on the circle: cardinal L-splines revisited. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1459–1473, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fine:1980:CSP

[119] Benjamin Fine. Congruence subgroups of the Picard group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1474–1481, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reisner:1980:OWF

[120] Shlomo Reisner. Operators which factor through convex Banach lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 1482–1500, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahman:1980:PFN

[121] Mizan Rahman. A product formula and a nonnegative Poisson kernel for Racah– Wilson polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1501–1517, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Banaschewski:1980:ISA

[122] B. Banaschewski. Injective sheaves of abelian groups: a counterexample. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):1518–1521, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dobbs:1980:CCI

[123] David E. Dobbs. Corrigendum: “On characterizing injective sheaves”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??): 1522, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [2].

Orzech:1980:CBG

[124] Morris Orzech. Correction to: “On the Brauer group of algebras having a grading and an action”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):1523–1524, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [1].

Wormald:1981:NPM

[125] Nicholas C. Wormald. On the number of planar maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1–11, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jackson:1981:GTR

[126] D. M. Jackson and I. P. Goulden. The generalisation of Tutte’s result for chromatic trees, by Lagrangian methods. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):12–19, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Oxley:1981:CM

[127] James G. Oxley. On 3-connected matroids. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):20–27, ????

1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Poguntke:1981:FGL

[128] W. Poguntke and B. Sands. On finitely generated lattices of finite width. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 28–48, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thomas:1981:NYR

[129] Glânffrwd P. Thomas. A note on Young’s raising operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):49–54, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Takeuchi:1981:CNM

[130] Hiroshi Takeuchi. On the class number of the maximal real subfield of a cyclotomic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):55–58, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fleming:1981:I

[131] R. J. Fleming and J. E. Jamison.

Isometries of s_{p}(α). Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):59–67, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Takahasi:1981:CQB

[132] Sin ei Takahasi. On the center of quasicentral Banach algebras with bounded approximate identity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):68–90, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartley:1981:KFP

[133] Richard Hartley. Knots with free period. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):91–102, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Huckaba:1981:L

[134] James A. Huckaba and Ira J. Papick. A localization of R[x]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):103–115, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ishibashi:1981:GOG

[135] Hiroyuki Ishibashi. Generators of orthogonal groups over valuation rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):116–128, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hanlon:1981:AAT

[136] Phil Hanlon. Algebras of acyclic type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):129–141, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Espelie:1981:SPC

[137] M. Solveig Espelie and James E. Joseph. Some properties of heta-closure. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):

142–149, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Daverman:1981:SUI

[138] R. J. Daverman and L. D. Loveland.

Any 2-sphere in E^{3} with uniform interior tangent balls is flat. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):150–167, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bass:1981:TVD

[139] D. W. Bass. Topologically versal deformations of matrices; codimension at most two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):168–180, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kulkarni:1981:GPR

[140] S. H. Kulkarni and B. V. Limaye. Gleason parts of real function algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):181–200, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1981:MBI

[141] Peter B. Borwein. Markov’s and Bernstein’s inequalities on disjoint intervals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):201–209, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Binding:1981:VMO

[142] Paul Binding. Variational methods for one and several parameter nonlinear eigenvalue problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):210–228, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ahlbrandt:1981:NSC

[143] Calvin D. Ahlbrandt, Don B. Hinton, and Roger T. Lewis. Necessary and sufficient conditions for the discreteness of the spectrum of certain singular differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):229–246, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Powell:1981:TTD

[144] Charles H. Powell and William R. Wade. Term by term dyadic differentiation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):247–256, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Furuta:1981:ECO

[145] Takayuki Furuta. Essentially convexoid operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):257–274, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Liang:1981:AHS

[146] Chao Chu Liang. σ_{4}-actions on homotopy spheres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):275–281, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Taylor:1981:DPI

[147] Alan D. Taylor. Diamond principles, ideals and the normal Moore space problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):282–296, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martin:1981:AAR

[148] John R. Martin. Absolute approximate retracts and AR-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):297–301, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jardine:1981:AHT

[149] J. F. Jardine. Algebraic homotopy theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):302–319, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Esterle:1981:IRD

[150] J. Esterle. Irregularity of the rate of decrease of sequences of powers in the Volterra algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):320–324, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heinicke:1981:LCP

[151] A. G. Heinicke. Localization and completion at primes generated by normalizing sequences in right Noetherian rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):325–346, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dickson:1981:EAA

[152] D. G. Dickson. Expansions of arbitrary analytic functions in series of exponentials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):347–356, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Egghe:1981:SCP

[153] Leo Egghe. Strong convergence of pramarts in Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):357–361, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Riemenschneider:1981:BIR

[154] S. D. Riemenschneider and A. Sharma. Birkhoff interpolation at the nth roots of unity: convergence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):362–371, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hegde:1981:WCR

[155] Shankar Hegde. Wallman compactification and representation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):372–380, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gordon:1981:SCL

[156] C. McA. Gordon, R. A. Litherland, and K. Murasugi. Signatures of covering links. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):381–394, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Walker:1981:RRF

[157] W. J. Walker. Rankings and ranking functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):395–399, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heinzer:1981:EID

[158] William Heinzer. An essential integral domain with a nonessential localization. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):400–403, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gratzer:1981:SFP

[159] G. Grätzer, A. P. Huhn, and H. Lakser.

On the structure of finitely presented lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):404–411, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pettet:1981:NPI

[160] Martin R. Pettet. Nilpotent partitioninducing automorphism groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 412–420, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Spackman:1981:NDS

[161] Kenneth W. Spackman. On the number and distribution of simultaneous solutions to diagonal congruences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):

421–436, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bannai:1981:UCS

[162] Eiichi Bannai and N. J. A. Sloane. Uniqueness of certain spherical codes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):437–449, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rogers:1981:CRM

[163] James T. Rogers, Jr. Completely regular mappings and homogeneous, aposyndetic continua. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):450–453, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dobbs:1981:LPC

[164] David E. Dobbs. Lying-over pairs of commutative rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):454–475, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bollobas:1981:CCG

[165] Béla Bollobás. Counting coloured graphs of high connectivity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 476–484, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kudla:1981:PDG

[166] Stephen S. Kudla and John J. Millson. The Poincaré dual of a geodesic algebraic curve in a quotient of the 2ball. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):485–499, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kestenband:1981:HCO

[167] Barbu C. Kestenband. Hermitian configurations in odd-dimensional projective geometries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):500–512, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shelstad:1981:EG

[168] D. Shelstad. Embeddings of L-groups.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):513–558, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jenkins:1981:EBL

[169] James A. Jenkins. On explicit bounds in Landau’s theorem. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):559–562, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hindman:1981:CEF

[170] Neil Hindman. On a conjecture of Erdős, Faber, and Lovász about n-colorings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):563–570, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohen:1981:LBT

[171] Jo-Ann D. Cohen. Locally bounded topologies on the ring of integers of a global field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):571–584, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Izawa:1981:MQR

[172] Tatsuo Izawa. Maximal quotient rings of endomorphism rings of E(R_{R})- torsionfree generators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):585–605, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leeming:1981:SPG

[173] D. J. Leeming and R. A. MacLeod. Some properties of generalized Euler numbers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):606–617, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Klee:1981:DRG

[174] Victor Klee and David Larman. Diameters of random graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):618–640, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heldermann:1981:DSN

[175] N. C. Heldermann. Developability and some new regularity axioms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 641–663, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Khan:1981:PRP

[176] M. A. Khan. Power-rich and powerdeficient LCA groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):664–670, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wilson:1981:IDM

[177] Leslie C. Wilson. Infinitely determined map germs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):671–684, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davidson:1981:CPR

[178] Kenneth R. Davidson. Compact perturbations of reflexive algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 685–700, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pierre:1981:PTA

[179] R. Pierre and Q. I. Rahman. On a problem of Turán about polynomials. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):701–733, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muller:1981:QPN

[180] Bruno J. Müller. The quotient problem for Noetherian rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):734–748, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Akcoglu:1981:DDA

[181] M. A. Akcoglu and A. del Junco. Differentiation of n-dimensional additive processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):749–768, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Balser:1981:LFS

[182] W. Balser, W. B. Jurkat, and A. Peyerimhoff. On linear functionals and summability factors for strong summability. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):769–781, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hansen:1981:SME

[183] Vagn Lundsgaard Hansen. Spaces of maps into Eilenberg–MacLane spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):782–785, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Khalil:1981:AM

[184] Roshdi Khalil. On the algebra of multipliers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):786–794, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Avis:1981:HSH

[185] David Avis. Hypermetric spaces and the Hamming cone. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):795–802, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wimp:1981:USE

[186] Jet Wimp. On uniqueness sets for expansions in sequences of functions arising from singular generating functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):803–816, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hill:1981:CTP

[187] Paul Hill. Criteria for total projectivity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):817–825, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hadwin:1981:DHD

[188] D. W. Hadwin. Dilations and Hahn decompositions for linear maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 826–839, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heerema:1981:DIA

[189] N. Heerema and T. Morrison. On derivations induced by p-adic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 840–856, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Williams:1981:CHD

[190] Scott W. Williams. Coabsolutes with homeomorphic dense subspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):

857–861, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Akemann:1981:UND

[191] Charles A. Akemann and Martin E. Walter. Unbounded negative definite functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):862–871, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vermeer:1981:PEC

[192] Hans Vermeer and Evert Wattel. Projective elements in categories with perfect heta-continuous maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):872–884, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edge:1981:SNQ

[193] W. L. Edge. A specialised net of quadrics having self-polar polyhedra, with details of the five-dimensional example. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):885–892, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gerhard:1981:WPO

[194] J. A. Gerhard and Mario Petrich. The word problem for orthogroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 893–900, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Musson:1981:IMP

[195] I. M. Musson. Irreducible modules for polycyclic group algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 901–914, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahman:1981:NRL

[196] Mizan Rahman. A nonnegative representation of the linearization coefficients of the product of Jacobi polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):915–928, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lowen-Colebunders:1981:OPM

[197] Eva Lowen-Colebunders. Open and proper maps characterized by continuous set-valued maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):929–936, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levesque:1981:SFU

[198] Claude Levesque. Systèmes fondamentaux d’unités de certains composés de deux corps quadratiques. I. (French) [Fundamental systems of units of certain compounds of two quadratic fields. I]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):937–945, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1981:GHM

[199] David Borwein and Amnon Jakimovski.

Generalization of the Hausdorff moment problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):946–960, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahman:1981:LPC

[200] Mizan Rahman. The linearization of the product of continuous q-jacobi polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):961–987, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorimer:1981:LCH

[201] J. W. Lorimer. Locally compact Hjelmslev planes and rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):988–1022, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cori:1981:PMW

[202] Robert Cori and Bernard Vauquelin. Planar maps are well labeled trees. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):1023–1042, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wolfe:1981:LPA

[203] Warren Wolfe. Limits on pairwise amicable orthogonal designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 1043–1054, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brillhart:1981:ITC

[204] John Brillhart, Michael Filaseta, and Andrew Odlyzko. On an irreducibility theorem of A. Cohn. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1055–1059, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1981:TPD

[205] N. L. Johnson. The translation planes of Dempwolff. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1060–1073, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mollin:1981:CDA

[206] R. A. Mollin. Cyclotomic division algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1074–1084, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deveney:1981:DSI

[207] James K. Deveney and John N. Mordeson. Distinguished subfields of intermediate fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1085–1096, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Huckleberry:1981:CHS

[208] A. T. Huckleberry and E. L. Livorni. A classification of homogeneous surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):1097–1110, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elliott:1981:GRE

[209] Joanne Elliott. Generalized resolvent equations and unsymmetric Dirichlet spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1111–1141, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kuiken:1981:MGR

[210] Kathryn Kuiken. On the monodromy groups of Riemann surfaces of genus geq1. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1142–1156, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kolaski:1981:IBS

[211] Clinton J. Kolaski. Isometries of Bergman spaces over bounded Runge domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1157–1164, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pellaumail:1981:WSS

[212] J. Pellaumail. Weak solutions for semimartingales. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1165–1181, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morita:1981:STC

[213] Jun Morita. On some twisted Chevalley groups over Laurent polynomial rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):1182–1201, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brouwer:1981:SMT

[214] A. E. Brouwer. On the size of a maximum transversal in a Steiner triple system. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1202–1204, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fialkow:1981:EST

[215] Lawrence A. Fialkow. Elements of spectral theory for generalized derivations. II. the semi-Fredholm domain. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 1205–1231, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ishibashi:1981:GQS

[216] Hiroyuki Ishibashi. Generators of U_{n}(V)

over a quasisemilocal semihereditary ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1232–1244, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pouzet:1981:WOS

[217] Maurice Pouzet and Ivan Rival. Which ordered sets have a complete linear extension? Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1245–1254, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Metzger:1981:BRS

[218] Thomas A. Metzger. On BMOA for Riemann surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1255–1260, ????

1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gasper:1981:OCF

[219] George Gasper. Orthogonality of certain functions with respect to complex valued weights. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1261–1270, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miers:1981:CLI

[220] C. Robert Miers. Closed Lie ideals in operator algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1271–1278, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kierstead:1981:RCH

[221] Henry A. Kierstead. Recursive colorings of highly recursive graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 1279–1290, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Radjabalipour:1981:RA

[222] Mehdi Radjabalipour. On reflexivity of algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1291–1308, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reilly:1981:SLV

[223] Norman R. Reilly. A subsemilattice of the lattice of varieties of lattice ordered groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1309–1318, ????

1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Katavolos:1981:NSR

[224] A. Katavolos. Are noncommutative L_{p} spaces really noncommutative? Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 1319–1327, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Savoyant:1981:RCA

[225] Michel Savoyant. Une remarque sur une classe d’algèbres de Dirichlet faibles ∗ de fonctions analytiques. (French) [A note on a class of weak ^{∗} Dirichlet algebras of analytic functions]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(6):1328– 1330, December 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ishibashi:1981:SPR

[226] Yasunori Ishibashi. Symbolic powers of regular primes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1331–1337, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ware:1981:VRR

[227] Roger Ware. Valuation rings and rigid elements in fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1338–1355, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Squires:1981:NCU

[228] W. A. Squires. Necessary conditions for universal interpolation in hatE′. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 1356–1364, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Phelps:1981:DPD

[229] K. T. Phelps. Direct product of derived Steiner systems using inversive planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):1365–1369, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mollin:1981:IQA

[230] Richard A. Mollin. Induced quaternion algebras in the Schur group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 1370–1379, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1981:SCL

[231] Kenneth W. Smith. Stability and categoricity of lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1380–1419, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brandenburg:1981:SCS

[232] Harald Brandenburg. Separating closed sets by continuous mappings into developable spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1420–1431, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yoshizawa:1981:BSS

[233] Mitsuo Yoshizawa. On block-schematic Steiner systems S(t, t + 1, v). Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 1432–1438, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kofner:1981:TOB

[234] Jacob Kofner. Transitivity and ortho-bases. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1439–1447, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Felix:1981:MBP

[235] Yves Félix. Modèles bifiltrés: une plaque tournante en homotopie rationnelle. (French) [Bifiltered models: a hub in rational homotopy]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 1448–1458, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chouinard:1981:KSD

[236] Leo G. Chouinard II. Krull semigroups and divisor class groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??): 1459–1468, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cecchini:1981:SRC

[237] C. Cecchini and A. Zappa. Some results on the center of an algebra of operators on VN(G) for the Heisenberg group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):1469–1486, ???? 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jha:1981:SAP

[238] Vikram Jha and Michael J. Kallaher. On spreads admitting projective linear groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):1487–1497, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bachman:1981:CSS

[239] George Bachman and Panagiotis D. Stratigos. Criteria for σ-smoothness, τsmoothness, and tightness of lattice regular measures, with applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(6):1498– 1525, December 1981. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sawyer:1982:HFP

[240] Eric Sawyer. Holomorphic functions with positive real part. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1–7, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Krempa:1982:IGR

[241] Jan Krempa. Isomorphic group rings of free abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):8–16, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Misso:1982:CCR

[242] Paola Misso. Commutativity conditions on rings with involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 17–22, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berberian:1982:FCR

[243] S. K. Berberian. Factor-correspondences in regular rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):23–30, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Halperin:1982:REF

[244] Stephen Halperin and Jean-Claude Thomas. Rational equivalence of fibrations with fibre G/K. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):31–43, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fournier:1982:HPN

[245] Gilles Fournier and Reine Fournier. On the homotopy property of Nussbaum’s fixed point index. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):44–62, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moody:1982:RSC

[246] R. V. Moody and T. Yokonuma. Root systems and Cartan matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 63–79, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bankston:1982:SOD

[247] Paul Bankston. Some obstacles to duality in topological algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 80–90, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [471].

Morris:1982:ESRa

[248] L. E. Morris. Eisenstein series for reductive groups over global function fields. I. the cusp form case. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):91–168, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roberts:1982:OSD

[249] Leslie G. Roberts. Ordinary singularities with decreasing Hilbert function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 169–180, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heitmann:1982:DD

[250] Raymond C. Heitmann and Stephen McAdam. Distinguished domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):181–193, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kirdar:1982:IRP

[251] M. S. Kirdar and T. H. R. Skyrme. On an identity relating to partitions and repetitions of parts. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):194–195, ????

1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1982:DPG

[252] D. D. Anderson and David F. Anderson. Divisibility properties of graded domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):196–215, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gauthier:1982:AHA

[253] P. M. Gauthier, W. Hengartner, and M. Labrèche. Approximation harmonique, approximation holomorphe et topologie. (French) [Harmonic approximation, holomorphic approximation, and topology]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):216–219, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baggett:1982:ABI

[254] Larry Baggett and Keith F. Taylor. On asymptotic behavior of induced representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):220–232, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ritter:1982:IGR

[255] Jürgen Ritter and Sudarshan Sehgal.

Integral group rings of some p-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):233–246, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yucas:1982:LLQ

[256] Joseph Yucas. Localizations of linked quaternionic mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):247–256, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jungnickel:1982:AGD

[257] Dieter Jungnickel. On automorphism groups of divisible designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 257–297, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pathak:1982:LIS

[258] V. D. Pathak. Linear isometries of spaces of absolutely continuous functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):298–306, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alspach:1982:CVT

[259] Brian Alspach and T. D. Parsons. A construction for vertex-transitive graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):307–318, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlson:1982:RPC

[260] Stephan C. Carlson. Rich proximities and compactifications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):319–348, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

John:1982:SLL

[261] David John. Simple links in locally compact connected Hausdorff spaces are nondegenerate. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):349–355, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herstein:1982:TDP

[262] I. N. Herstein. A theorem on derivations of prime rings with involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 356–369, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pedersen:1982:TJA

[263] Gert Kjærgård Pedersen and Erling Størmer. Traces on Jordan algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):370–373, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Akin:1982:EFG

[264] Ethan Akin. Exponential families and game dynamics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):374–405, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deeb:1982:NSC

[265] Waleed Deeb. Necessary and sufficient conditions for the equality of L(f) and l^{1}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):406–410, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Batola:1982:GFM

[266] François Batola. Une généralisation d’une formule de Meixner–Tricomi. (French) [A generalization of a formula of Meixner–Tricomi]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):411–422, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoo:1982:PIS

[267] C. S. Hoo. Pseudocomplemented and implicative semilattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):423–437, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anstee:1982:PCM

[268] R. P. Anstee. Properties of a class of (0, 1)-matrices covering a given matrix. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):438–453, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kochman:1982:PGA

[269] Stanley O. Kochman. Primitive generators for algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):454–465, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Axler:1982:TOB

[270] Sheldon Axler, John B. Conway, and Gerard McDonald. Toeplitz operators on Bergman spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):466–483, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Malzan:1982:PPR

[271] J. Malzan. Products of positive reflections in the orthogonal group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 484–499, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burris:1982:TUR

[272] Stanley Burris and John Lawrence.

Two undecidability results using modified Boolean powers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):500–505, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cross:1982:AII

[273] G. E. Cross. Additivity of the P^{n}- integral, II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):506–512, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goulden:1982:ALM

[274] I. P. Goulden and D. M. Jackson.

The application of Lagrangian methods to the enumeration of labelled trees with respect to edge partition. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 513–518, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Poljak:1982:PAC

[275] Svatopluk Poljak and Daniel Turzík. A polynomial algorithm for constructing a large bipartite subgraph, with an application to a satisfiability problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 519–524, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1982:TTP

[276] Jon L. Johnson. Three topological properties from Noetherian rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 525–534, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruns:1982:SFC

[277] Günter Bruns and Richard Greechie. Some finiteness conditions for orthomodular lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):535–549, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benkart:1982:RFD

[278] Georgia M. Benkart, Daniel J. Britten, and J. Marshall Osborn. Real flexible division algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):550–588, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ramras:1982:BBN

[279] Mark Ramras. Bounds on Betti numbers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):589–592, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dubuc:1982:REA

[280] Serge Dubuc and Jean Savoie. Résolution d’équations associées à un système de Tchebycheff. (French) [Solving equations associated with a Chebyshev system]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):593–620, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ball:1982:GOS

[281] Richard N. Ball. The generalized orthocompletion and strongly projectable hull of a lattice ordered group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 621–661, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goodaire:1982:CLW

[282] Edgar G. Goodaire and D. A. Robinson. A class of loops which are isomorphic to all loop isotopes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):662–672, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kumagai:1982:MSU

[283] Donna Kumagai. The measure spectrum of a uniform algebra and subharmonicity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):673–685, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Childs:1982:SCC

[284] L. N. Childs. A Stickelberger condition on cyclic Galois extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 686–690, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levy:1982:SRW

[285] Lawrence S. Levy and Patrick F. Smith. Semiprime rings whose homomorphic images are serial. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):691–695, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruns:1982:OLW

[286] Günter Bruns and Richard Greechie. Orthomodular lattices which can be covered by finitely many blocks. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 696–699, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fiedorowicz:1982:HOR

[287] Z. Fiedorowicz and J. P. May. Homology operations revisited. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):700–717, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bustoz:1982:AUP

[288] Joaquin Bustoz and Mourad E. H. Ismail. The associated ultraspherical polynomials and their q-analogues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 718–736, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aziz:1982:IPP

[289] Abdul Aziz. Inequalities for polynomials with a prescribed zero. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):737–740, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1982:CS

[290] W. T. Tutte. Chromatic solutions.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):741–758, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yoneda:1982:PSU

[291] Kaoru Yoneda. Perfect sets of uniqueness on the group 2^{ω}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):759–764, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorentz:1982:ACB

[292] G. G. Lorentz. The analytic character of the Birkhoff interpolation polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):765–768, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shastri:1982:DOM

[293] Anant R. Shastri. Degree one maps and a realization theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):769–796, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hill:1982:RWI

[294] David A. Hill. Rings whose indecomposable injective modules are uniserial. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):797–805, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beardon:1982:FDR

[295] A. F. Beardon and P. J. Nicholls. Ford and Dirichlet regions for Fuchsian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):806–815, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rigby:1982:HTI

[296] J. F. Rigby. Half-turns and infinite chains of Clifford configurations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 816–831, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zaheer:1982:LSV

[297] Neyamat Zaheer. On Lucas-sets for vector-valued abstract polynomials in K-inner product spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):832–852, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dauge:1982:OSE

[298] Monique Dauge. Opérateur de Stokes dans des espaces de Sobolev à poids sur des domaines anguleux. (French) [Stokes operator in Sobolev spaces with weights in angular regions]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):853–882, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lubin:1982:SIC

[299] A. R. Lubin. Spectral inclusion and c.n.e. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):883–887, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pierre:1982:PTA

[300] R. Pierre and Q. I. Rahman. On a problem of Turán about polynomials. III. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):888–899, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pastijn:1982:LCD

[301] Francis Pastijn. L’enveloppe conjuguée d’un demi-groupe inverse à métacentre idempotent. (French) [The conjugated envelope of an inverse semigroup of idempotent metacenter]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):900–909, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kolaski:1982:IWB

[302] Clinton J. Kolaski. Isometries of weighted Bergman spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 910–915, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnston:1982:RPR

[303] Elgin H. Johnston. Rearrangements that preserve rates of divergence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 916–920, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Katchalski:1982:IPF

[304] M. Katchalski and A. Liu. Intersection patterns of families of convex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 921–931, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frappier:1982:IB

[305] C. Frappier and Q. I. Rahman. On an inequality of S. Bernstein. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):932–944, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chang:1982:EOC

[306] Bomshik Chang. Elements of order Coxeter number +1 in Chevalley groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):945–951, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1982:CSI

[307] W. T. Tutte. Chromatic solutions. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):952–960, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Newman:1982:PCM

[308] D. J. Newman and T. J. Rivlin. On polynomials with curved majorants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 961–968, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Williams:1982:CNB

[309] Kenneth S. Williams and James D. Currie. Class numbers and biquadratic reciprocity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):969–988, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Meril:1982:SH

[310] A. Meril. Symbole holomorphe. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 989–995, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:1982:BGC

[311] Heisook Lee and Morris Orzech. Brauer groups, class groups and maximal orders for a Krull scheme. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):996–1010, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Evans:1982:CEC

[312] Ronald Evans, Mourad E. H. Ismail, and Dennis Stanton. Coefficients in expansions of certain rational functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):

1011–1024, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rosenoer:1982:CRO

[313] Shlomo Rosenoer. Completely reducible operator algebras and spectral synthesis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):1025–1035, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schmerl:1982:EVB

[314] James H. Schmerl. The effective version of Brooks’ theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1036–1046, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Argyros:1982:PSS

[315] S. Argyros and N. Kalamidas. Property k_{α,n} on spaces with strictly positive measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1047–1058, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levesque:1982:SFU

[316] Claude Levesque. Systèmes fondamentaux d’unités de certains corps de degré 4 et de degré 8 sur Q. (French) [Fundamental systems of units of some objects of degree 4 and degree 8 on Q]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 1059–1090, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Watson:1982:LCN

[317] W. Stephen Watson. Locally compact normal spaces in the constructible universe. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1091–1096, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wolf:1982:SNS

[318] Thomas R. Wolf. Solvable and nilpotent subgroups of GL(n, q^{m}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 1097–1111, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morris:1982:ESRb

[319] L. E. Morris. Eisenstein series for reductive groups over global function fields. II. the general case. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1112–1182, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fegan:1982:SFP

[320] H. D. Fegan. Special function potentials for the Laplacian. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1183–1194, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kantor:1982:OTP

[321] William M. Kantor. Ovoids and translation planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1195–1207, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Katavolos:1982:NSI

[322] A. Katavolos. Noncommutative L_{p} spaces. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1208–1214, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Santharoubane:1982:KML

[323] L. J. Santharoubane. Kac–Moody Lie algebras and the classification of nilpotent Lie algebras of maximal rank. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 1215–1239, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Matlis:1982:CCR

[324] Eben Matlis. Commutative coherent rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1240–1244, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanDaele:1982:DSL

[325] A. van Daele. Dense subalgebras of left Hilbert algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1245–1250, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hilton:1982:LSP

[326] A. J. W. Hilton and C. A. Rodger. Latin squares with prescribed diagonals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):1251–1254, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ohta:1982:TEP

[327] Haruto Ohta. Topological extension properties and projective covers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 1255–1275, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carson:1982:DWR

[328] Andrew B. Carson and Murray A. Marshall. Decomposition of Witt rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):1276–1302, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kieffer:1982:PET

[329] John C. Kieffer and Maurice Rahe. The pointwise ergodic theorem for transformations whose orbits contain or are contained in the orbits of a measurepreserving transformation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1303– 1318, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alexander:1982:NPC

[330] H. Alexander. A note on projective capacity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1319–1329, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kent:1982:CCL

[331] D. C. Kent and G. D. Richardson. A compactification with heta-continuous lifting property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1330–1334, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilbert:1982:CNT

[332] William J. Gilbert. Complex numbers with three radix expansions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 1335–1348, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coelho:1982:GRU

[333] Sônia P. Coelho. Group rings with units of bounded exponent over the center. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (??):1349–1364, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heath-Brown:1982:PNS

[334] D. R. Heath-Brown. Prime numbers in short intervals and a generalized Vaughan identity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1365–1377, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dow:1982:SSS

[335] Alan Dow. Some separable spaces and remote points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1378–1389, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Filaseta:1982:FGI

[336] Michael Filaseta. A further generalization of an irreducibility theorem of A. Cohn. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):1390–1395, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Birman:1982:ECP

[337] Joan S. Birman. Errata: “On the Conjugacy Problem in the Braid Group”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34 (6):1396–1397, December 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCool:1982:ERB

[338] James McCool. Erratum: “On reducible braids” [Word problems, II (Oxford, 1976), 261–295, North-Holland, Amsterdam, 1980. MR 81m:20058]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??): 1398, ???? 1982. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dangello:1983:SPC

[339] F. Dangello and R. Lindahl. Semidirect product compactifications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 1–32, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Enns:1983:RP

[340] T. C. Enns. The radii of polyhedrons.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 35 (??):33–42, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edwards:1983:SCC

[341] Robert Edwards and David Gillman. Any spine of the cube is 2-collapsible. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35 (??):43–48, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Paterson:1983:ATE

[342] Alan L. T. Paterson. Amenability and translation experiments. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 49–58, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gluck:1983:TSS

[343] David Gluck. Trivial set-stabilizers in finite permutation groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 59–67, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bisztriczky:1983:NSO

[344] Tibor Bisztriczky. C-nodal surfaces of order three. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):68–100, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Monson:1983:SQF

[345] Barry Monson. Simplicial quadratic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):101–116, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barnes:1983:NLP

[346] E. S. Barnes and N. J. A. Sloane.

New lattice packings of spheres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 117–130, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1983:REI

[347] P. F. Smith. Rings with enough invertible ideals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):131–144, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baxter:1983:WSE

[348] J. R. Baxter and J. H. Olsen. Weighted and subsequential ergodic theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35 (??):145–166, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kirschenbaum:1983:SWJ

[349] Marc Kirschenbaum and Yung Chen Lu.

Sufficiency of Weierstrass jets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 167–176, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoffman:1983:HAB

[350] P. Hoffman. Hopf algebras from branching rules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):177–192, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hurvitz:1983:IML

[351] Sara Hurvitz. l′-isolated maps and localizations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):193–217, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fresnel:1983:PTT

[352] Jean Fresnel and Michel Matignon. Produit tensoriel topologique de corps valués. (French) [Topological tensor product of valued objects]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 218–273, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fong:1983:UIO

[353] C. K. Fong and J. A. R. Holbrook. Unitarily invariant operator norms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 274–299, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bergen:1983:DIV

[354] Jeffrey Bergen, I. N. Herstein, and Charles Lanski. Derivations with invertible values. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):300–310, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Van:1983:RPT

[355] Quach thi Câm Vân. On a realization of prime tangles and knots. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):

311–323, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beesack:1983:ICJ

[356] Paul R. Beesack. On inequalities complementary to Jensen’s. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):324–338, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Friedman:1983:MS

[357] Nathaniel A. Friedman. Mixing on sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):339–352, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pavlakos:1983:IPO

[358] Panaiotis K. Pavlakos. On integration in partially ordered groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 353–372, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Takahashi:1983:CDC

[359] Sin ei Takahashi. Central double centralizers on quasicentral Banach algebras with bounded approximate identity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):373–384, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fantham:1983:GEC

[360] P. H. H. Fantham and E. J. Moore.

Groupoid enriched categories and homotopy theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):385–416, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Liskovets:1983:ENC

[361] V. A. Liskovets and T. R. S. Walsh. The enumeration of nonisomorphic 2connected planar maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):417–435, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edelson:1983:ABN

[362] Allan L. Edelson and Emilia Perri. Asymptotic behaviour of nonoscillatory equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):436–453, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schwartz:1983:LSS

[363] Niels Schwartz. Local stability and saturation in spaces of orderings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 454–477, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pandey:1983:HTG

[364] J. N. Pandey and Muhammad Aslam Chaudhry. The Hilbert transform of generalized functions and applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35 (??):478–495, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cenzer:1983:RWF

[365] Douglas Cenzer and R. Daniel Mauldin.

Representations of well-founded preference orders. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):496–508, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bray:1983:PCD

[366] Una Bray and George Whaples. Polynomials with coefficients from a division ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):509–515, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moore:1983:MFP

[367] R. L. Moore and G. Weiss. The metric Fuglede property and normality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 516–525, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leeming:1983:GEN

[368] D. J. Leeming and R. A. MacLeod.

Generalized Euler number sequences: asymptotic estimates and congruences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35 (??):526–546, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1983:UAP

[369] Peter B. Borwein. Uniform approximation by polynomials with variable exponents. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):547–557, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Babiker:1983:HCS

[370] A. G. A. G. Babiker and S. Graf. Homomorphism-compact spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 558–576, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kan:1983:EPL

[371] Charn-Huen Kan. Ergodic properties of Lamperti operators. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):577–588, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vonzurGathen:1983:SSC

[372] Joachim von zur Gathen. Secant spaces to curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):589–612, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chattopadhyay:1983:MSE

[373] K. C. Chattopadhyay, Olav Njåstad, and W. J. Thron. Merotopic spaces and extensions of closure spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 613–629, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ucci:1983:X

[374] Jack Ucci. On BP⟨1⟩^{∗}(K(Z, 3); Z/p). Canadian Journal of Mathematics = Journal canadien de mathématiques, 35 (4):630–653, August 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Adjaero:1983:UCR

[375] Isabelle Adjaero and Eugene Spiegel.

On the uniqueness of the coefficient ring in a group ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):654–673, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koblitz:1983:AEF

[376] Neal Koblitz. p-adic eigen-functions for Kubert distributions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):674–686, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gessel:1983:SCG

[377] Ira M. Gessel. Some congruences for generalized Euler numbers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):687–709, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miers:1983:AID

[378] C. Robert Miers and John Phillips.

Algebraic inner derivations on operator algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):710–723, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1983:BTD

[379] T. C. Brown and J. P. Buhler. Behrend’s theorem for dense subsets of finite vector spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):724–734, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ahern:1983:PIS

[380] Patrick Ahern. The Poisson integral of a singular measure. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):735–749, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lebow:1983:SA

[381] A. Lebow and M. Schreiber. Scalar actions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):750–768, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McNulty:1983:IPT

[382] George F. McNulty, T. Nordahl, and H. E. Scheiblich. Injectives and projectives in term finite varieties of algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):769–775, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herz:1983:AFE

[383] Carl Herz. Alternating 3-forms and exceptional simple Lie groups of type G_{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):776–806, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Daniels:1983:NLC

[384] Peg Daniels. Normal, locally compact, boundedly metacompact spaces are paracompact: an application of Pixley–Roy spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):807–823, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bell:1983:TBA

[385] Murray Bell. Two Boolean algebras with extreme cellular and compactness properties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):824–838, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Noussair:1983:GPS

[386] Ezzat S. Noussair and Charles A. Swanson. Global positive solutions of semilinear elliptic equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):839–861, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kannappan:1983:GFE

[387] Pl. Kannappan and C. T. Ng. On a generalized fundamental equation of information. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):862–872, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Browne:1983:DOT

[388] Patrick J. Browne and R. V. Nillsen. On difference operators and their factorization. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):873–897, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moody:1983:MBR

[389] R. V. Moody and A. Pianzola. lambdamappings between representation rings of Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):898–960, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kandil:1983:LID

[390] A. Kandil. On large inductive dimension of proximity spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):961–973, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morris:1983:SRE

[391] Lawrence Morris. Some remarks on Eisenstein series for metaplectic coverings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):974–985, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Terada:1983:DSP

[392] Toshiji Terada. Dense subspaces of product spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):986–1000, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Warren:1983:SUU

[393] Richard H. Warren. Spaces with a unique uniformity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):1001–1009, ???? 1983. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Huijsmans:1983:MIR

[394] Charles B. Huijsmans and Ben de Pagter.

Maximal d-ideals in a Riesz space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 1010–1029, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sherling:1983:CCH

[395] Dorothy D. Sherling. Concerning the cone = hyperspace property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 1030–1048, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yamada:1983:RDB

[396] Akio Yamada and Yukio Matsumoto.

On the resolution diagrams of the Brieskorn singularities (2, q, r) of type II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35 (??):1049–1058, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hanche-Olsen:1983:STP

[397] Harald Hanche-Olsen. On the structure and tensor products of JC-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35 (??):1059–1074, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shahidi:1983:SRI

[398] Freydoon Shahidi. Some results on Lindistinguishability for SL(r). Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??):

1075–1109, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Strelitz:1983:TTG

[399] Sh. Strelitz. Three theorems on the growth of entire transcendental solutions of algebraic differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(??): 1110–1128, ???? 1983. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

OBrien:1983:OPR

[400] G. L. O’Brien. Obtaining prescribed rates of convergence for the ergodic theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 35(6):1129–1146, December 1983. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boivin:1984:AHS

[401] A. Boivin and P. M. Gauthier. Approximation harmonique sur les surfaces de Riemann. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):1–8, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gervais:1984:DVS

[402] Jean-Jacques Gervais. Déformations Gverselles et G-stables. (French) [G-versal and G-stable deformations]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??): 9–21, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Singh:1984:SRN

[403] Surjeet Singh. Serial right Noetherian rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):22–37, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Swardson:1984:CCC

[404] Mary Anne Swardson. The character of certain closed sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):38–57, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lowen-Colebunders:1984:RKC

[405] Eva Lowen-Colebunders. On the regularity of the Kowalsky completion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36 (??):58–70, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Allegretto:1984:NSW

[406] Walter Allegretto. Nonnegative solutions for a weakly nonlinear elliptic equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):71–83, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Oudadess:1984:RRA

[407] Mohamed Oudadess. Rayon de regularité dans les algèbres infraséquentielles. (French) [Radius of regularity in infrasequential algebras]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):84–94, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1984:DLM

[408] J. M. Borwein and H. M. Strójwas. Directionally Lipschitzian mappings on Baire spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):95–130, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edge:1984:TSN

[409] W. L. Edge. Tangent spaces of a normal surface with hyperelliptic sections. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36 (??):131–143, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tholen:1984:PCM

[410] Walter Tholen. Pro-categories and multiadjoint functors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):144–155, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1984:NST

[411] Daniel A. Marcus. Normal semimodules: a theory of generalized convex cones. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36 (??):156–177, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Varadarajan:1984:PA

[412] K. Varadarajan. Projective approximations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):178–192, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Khuri:1984:ERG

[413] Soumaya Makdissi Khuri. Endomorphism rings and Gabriel topologies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36 (??):193–205, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jespers:1984:CKR

[414] E. Jespers and P. Wauters. On central Ω-krull rings and their class groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36 (??):206–239, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stout:1984:SBV

[415] Edgar Lee Stout. Smooth boundary values along totally real submanifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36 (??):240–248, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Small:1984:DEL

[416] Charles Small. Diagonal equations over large finite fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):249–262, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koubek:1984:TMB

[417] Václav Koubek. Towards minimal binding varieties of lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):263–285, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1984:HMM

[418] J. M. Borwein and B. Richmond. How many matrices have roots? Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(2): 286–299, April 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Djokovic:1984:PRG

[419] Dragomir Ž. – Doković and Jerry Malzan. Products of reflections in the group SO^{∗}(2n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):300–326, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pin:1984:SPR

[420] J.-E. Pin. Semigroupe des parties et relations de Green. (French) [Semigroup of parts and Green’s relations]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??): 327–343, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Iroz:1984:API

[421] Juana Iroz and David E. Rush. Associated prime ideals in non-Noetherian rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):344–360, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vanstone:1984:CDM

[422] J. R. Vanstone. A canonical decomposition in mixed exterior algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):

361–373, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruce:1984:DGR

[423] J. W. Bruce and C. G. Gibson. Distancegenericity for real algebraic hypersurfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):374–384, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ono:1984:IEE

[424] Akira Ono. Interior estimates for elliptic partial differential equations in the L^{(q,λ)} spaces of strong type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??): 385–404, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barnes:1984:SHI

[425] David C. Barnes. Supplements of Hölder’s inequality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):405–420, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Giambruno:1984:SHR

[426] A. Giambruno. On the symmetric hypercenter of a ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):421–435, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wright:1984:WCO

[427] Steve Wright. Weakly compact, operator-valued derivations of type I von Neumann algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):436–457, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Forelli:1984:ND

[428] Frank Forelli. A note on divisibility in H^{∞}(X). Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):458–469, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Knuppel:1984:CHP

[429] Frieder Knüppel. A class of homomorphisms of pre-Hjelmslev groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??): 470–494, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Adamek:1984:FPE

[430] Jirí Adámek and Wolfgang Merzenich. Fixed points as equations and solutions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36 (??):495–519, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tinsley:1984:VOO

[431] Fred Tinsley and David G. Wright. On vertical order of one-dimensional compacta in E^{3}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):520–528, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yuster:1984:GAT

[432] Thomas Yuster and Bruce Peterson. A generalization of an addition theorem for solvable groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):529–536, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bisztriczky:1984:ICS

[433] Tibor Bisztriczky. Inflectional convex space curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):537–549, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Higgs:1984:ITC

[434] Denis Higgs. Injectivity in the topos of complete Heyting algebra valued sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36 (3):550–568, June 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barany:1984:DCF

[435] Imre Bárány, Zoltán Füredi, and János Pach. Discrete convex functions and proof of the six circle conjecture of Fejes Tóth. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):569–576, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brook:1984:DS

[436] Cecilia H. Brook. Decompositions of submeasures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):577–590, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hilton:1984:SGA

[437] Peter Hilton. On special groupautomorphisms and their composition. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36 (??):591–600, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vauthier:1984:FGS

[438] J. Vauthier. Fonction de Green et surfaces nodales des fonctions propres de l’hamiltonien −\frac{1}{2}Δ+V. (French) [Green function and nodal surfaces of the eigenfunctions of the Hamiltonian −\frac{1}{2}Δ + V]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36 (??):601–614, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Geller:1984:SHW

[439] Daryl Geller. Spherical harmonics, the Weyl transform and the Fourier transform on the Heisenberg group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??): 615–684, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dunkl:1984:OPS

[440] Charles F. Dunkl. Orthogonal polynomials with symmetry of order three. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??): 685–717, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hahn:1984:ABN

[441] Kyong T. Hahn. Asymptotic behavior of normal mappings of several complex variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):718–746, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kobayashi:1984:RSB

[442] Shoji Kobayashi. Range sets and BMO norms of analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??): 747–755, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Laczkovich:1984:DPC

[443] M. Laczkovich. On the difference property of the class of pointwise discontinuous functions and of some related classes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):756–768, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jha:1984:TPW

[444] Vikram Jha. On translation planes which admit solvable autotopism groups having a large slope orbit. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??): 769–782, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kohli:1984:UVC

[445] J. K. Kohli. A unified view of (complete) regularity and certain variants of (complete) regularity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):783–794, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoo:1984:MAS

[446] C. S. Hoo and P. V. Ramana Murty. Modular and admissible semilattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36 (??):795–799, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ogg:1984:SSC

[447] A. P. Ogg. Some special classes of Cartan matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):800–819, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fillmore:1984:ILC

[448] P. A. Fillmore and W. E. Longstaff. On isomorphisms of lattices of closed subspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):820–829, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bion-Nadal:1984:EEN

[449] Jocelyne Bion-Nadal. Espace des états normaux d’un facteur de type III_{λ}; (0 < λ < 1) et d’un facteur de type III_{0}. (French) [Normal state space of a factor of type III_{λ}; (0 < λ < 1) and of a factor of type III_{0}]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):830–882, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jouak:1984:EFC

[450] Mohamed Jouak and Lionel Thibault. Equicontinuity of families of convex and concave-convex operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??): 883–898, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Assem:1984:TTI

[451] Ibrahim Assem. Torsion theories induced by tilting modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??): 899–913, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hughes:1984:UVO

[452] Rhonda J. Hughes. On unbounded Volterra operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):914–923, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wong:1984:GMC

[453] R. Wong and J. P. McClure. Generalized Mellin convolutions and their asymptotic expansions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):924–960, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Winter:1984:ODC

[454] David J. Winter. Outer derivations and classical-Albert–Zassenhaus Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):961–972, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cain:1984:SSP

[455] Bryan E. Cain. Some spectral properties of polar decompositions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(6): 973–985, December 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lemaire:1984:LRA

[456] Claude Lemaire. Localisation relative associée à une théorie de torsion. (French) [Relative localization associated with a theory of torsion]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??): 986–999, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Graves:1984:CWC

[457] William H. Graves and Wolfgang Ruess. Compactness and weak compactness in spaces of compact-range vector measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):1000–1020, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bures:1984:MLP

[458] Donald Bures. Modularity in the lattice of projections of a von Neumann algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):1021–1030, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pierre:1984:EDP

[459] R. Pierre. On explicit decomposition for positive polynomials on [−1, +1] with applications to extremal problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36 (??):1031–1045, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wong:1984:GFF

[460] H. Ship-Fah Wong. A general form of the functional LIL for Banach-valued Brownian motion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):1046–1066, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Meier:1984:TFG

[461] David Meier and Akbar Rhemtulla. On torsion-free groups of finite rank. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??): 1067–1080, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McAuley:1984:DOC

[462] Louis F. McAuley and Eric E. Robinson. Discrete open and closed mappings on generalized continua and Newman’s property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):1081–1112, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See correction [506].

Banaschewski:1984:MCH

[463] B. Banaschewski. More on compact Hausdorff spaces and finitary duality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36 (??):1113–1118, ???? 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fenn:1984:BCM

[464] Roger Fenn and Denis Sjerve. Basic commutators and minimal Massey products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(6):1119–1146, December 1984. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1985:GSJ

[465] Howard Smith. Groups with the subnormal join property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(1):1–16, February 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Flapan:1985:IPK

[466] Erica Flapan. Infinitely periodic knots. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37 (1):17–28, February 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bami:1985:RFS

[467] M. Lashkarizadeh Bami. Representations of foundation semigroups and their algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(1):29–47, February 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Silverman:1985:SSF

[468] H. Silverman and E. M. Silvia. Subclasses of starlike functions subordinate to convex functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(1):48–61, February 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Novinger:1985:LIS

[469] W. P. Novinger and D. M. Oberlin. Linear isometries of some normed spaces of analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(1):62–74, February 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mathieu:1985:RNF

[470] Philippe Mathieu. Résultats nouveaux sur les faisceaux quasi-unipotents. (French) [New results on quasiunipotent bundles]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(1):75–81, February 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bankston:1985:CSO

[471] Paul Bankston. Corrigendum to: “Some obstacles to duality in topological algebra”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(1):82–83, February 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [247].

Pathak:1985:HTS

[472] R. S. Pathak. On Hankel transformable spaces and a Cauchy problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(1): 84–106, February 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Granirer:1985:GTP

[473] Edmond E. Granirer. Geometric and topological properties of certain w^{∗} compact convex sets which arise from the study of invariant means. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(1): 107–121, February 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Britten:1985:BC

[474] D. J. Britten and F. W. Lemire. On basic cycles of A_{n}, B_{n}, C_{n} and D_{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(1): 122–140, February 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Andrews:1985:UPI

[475] Kevin T. Andrews. Universal Pettis integrability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):141–159, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bratteli:1985:DTC

[476] Ola Bratteli and Frederick M. Goodman. Derivations tangential to compact group actions: spectral conditions in the weak closure. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):160–192, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Skandalis:1985:ESK

[477] Georges Skandalis. Exact sequences for the Kasparov groups of graded algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):193–216, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bisztriczky:1985:VTC

[478] Tibor Bisztriczky. An n-vertex theorem for convex space curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):217–237, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dixon:1985:OSP

[479] John D. Dixon. The orbit-stabilizer problem for linear groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??): 238–259, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frangos:1985:CVV

[480] Nikos E. Frangos. On convergence of vector valued pramarts and subpramarts. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):260–270, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gerhard:1985:FIC

[481] J. A. Gerhard and Mario Petrich.

Free involutorial completely simple semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):271–295, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

James:1985:CPS

[482] Donald G. James. Collineations of polar spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):296–309, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Essen:1985:ODO

[483] M. Essén. Optimization and αdisfocality for ordinary differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):310–323, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gratzer:1985:TGA

[484] George Grätzer and David Kelly. A technique to generate m-ary free lattices from finitary ones. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):324–336, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Williams:1985:CBR

[485] Kenneth S. Williams, Christian Friesen, and Lawrence J. Howe. Criteria for biquadratic residuacity modulo a prime p involving quaternary representations of p. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37 (??):337–370, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moricz:1985:RCS

[486] Ferenc Móricz. On the restricted Cesàro summability of multiple orthogonal series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):371–384, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Comez:1985:DMS

[487] Dogan Çömez. Differentiation of multiparameter superadditive processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37 (??):385–404, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Solel:1985:AAO

[488] Baruch Solel. Algebras of analytic operators associated with a periodic flow on a von Neumann algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):405–429, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cac:1985:MUL

[489] Nguyên Phuong Các. The method of upper and lower solutions for some nonlinear boundary value problems in unbounded domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):430–441, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gluck:1985:LIC

[490] David Gluck. The largest irreducible character degree of a finite group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??): 442–451, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jain:1985:HR

[491] S. K. Jain and D. S. Malik. q-hypercyclic rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):452–466, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gordon:1985:NRH

[492] Carolyn S. Gordon. Naturally reductive homogeneous Riemannian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):467–487, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Soule:1985:OTA

[493] Christophe Soulé. Opérations en Kthéorie algébrique. (French) [Operations in algebraic K-theory]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):488–550, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahman:1985:EFB

[494] Mizan Rahman. A q-extension of Feldheim’s bilinear sum for Jacobi polynomials and some applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??): 551–578, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ho:1985:FPP

[495] Chat Yin Ho. Finite projective planes that admit a strongly irreducible collineation group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):579–611, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Comez:1985:ETM

[496] Dogan Çömez. An ergodic theorem for multidimensional superadditive processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):612–634, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Holzherr:1985:GFD

[497] A. K. Holzherr. Groups with finitedimensional irreducible multiplier representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):635–643, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levesque:1985:ISU

[498] Claude Levesque. An independent system of units in certain algebraic number fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):644–663, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Magyar:1985:DAI

[499] Zoltán Magyar and Zoltán Sebestyén.

On the definition of C^{∗}-algebras. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37 (??):664–681, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grzaslewicz:1985:EPC

[500] Ryszard Grzaślewicz. Extreme positive contractions on finite-dimensional lspp-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):682–699, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hussein:1985:GS

[501] D. Hussein, M. A. Natsheh, and I. Qumsiyeh. Generalized L(f) spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??): 700–709, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Flath:1985:BEA

[502] Daniel E. Flath and L. C. Biedenharn. Beyond the enveloping algebra of sl_{3}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37 (??):710–729, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steprans:1985:SSV

[503] Juris Steprans. Strong Q-sequences and variations on Martin’s axiom. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??): 730–746, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Curtis:1985:AST

[504] D. W. Curtis. Application of a selection theorem to hyperspace contractibility. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37 (??):747–759, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gaveau:1985:DPS

[505] Bernard Gaveau and Jacques Vauthier. The Dirichlet problem for the subelliptic Laplacian on the Heisenberg group. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37 (??):760–766, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McAuley:1985:CDO

[506] L. F. McAuley and E. E. Robinson. A correction to: “Discrete open and closed mappings on generalized continua and Newman’s property”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):767–768, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [462].

Quigg:1985:APF

[507] John C. Quigg. Approximately periodic functionals on C^{∗}-algebras and von Neumann algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):769–784, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bami:1985:BTH

[508] M. Lashkarizadeh Bami. Bochner’s theorem and the Hausdorff moment theorem on foundation topological semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37 (??):785–809, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoffmann:1985:IHC

[509] Rudolf-E. Hoffmann. The injective hull and the scrCscrL-compactification of a continuous poset. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):810–853, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bender:1985:AAC

[510] Edward A. Bender and Nicholas C.

Wormald. Almost all convex polyhedra are asymmetric. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):854–871, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kraft:1985:SMC

[511] Jürgen Kraft. Singularity of monomial curves in A^{3} and Gorenstein monomial curves in A^{4}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):872–892, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ng:1985:NMC

[512] Tze Beng Ng. A note on the mod 2 cohomology of BhatSO_{n}⟨16⟩. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??): 893–907, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Andrew:1985:PTL

[513] A. D. Andrew. Projections on treelike Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):908–920, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schep:1985:CPC

[514] Anton R. Schep. Compactness properties of Carleman and Hille–Tamarkin operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):921–933, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Parks:1985:NSG

[515] Alan E. Parks. Nilpotent by supersolvable M-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):934–962, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bankston:1985:MFA

[516] Paul Bankston and Richard Schutt. On minimally free algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):963–978, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frenzen:1985:UAE

[517] C. L. Frenzen and R. Wong. A uniform asymptotic expansion of the Jacobi polynomials with error bounds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??): 979–1007, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Katre:1985:ECQ

[518] S. A. Katre and A. R. Rajwade. Euler’s criterion for quintic nonresidues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??): 1008–1024, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wigley:1985:CBS

[519] Neil M. Wigley. Corner behavior of solutions of semilinear Dirichlet problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):1025–1046, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beattie:1985:AAA

[520] Margaret Beattie. Automorphisms of G-azumaya algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):1047–1058, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shum:1985:CPS

[521] K. P. Shum, M. W. Chan, C. K. Lai, and K. Y. So. Characterizations for prime semilattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):1059–1073, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Treiman:1985:GGL

[522] Jay S. Treiman. Generalized gradients, Lipschitz behavior and directional derivatives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):1074–1084, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goodman:1985:ASC

[523] T. N. T. Goodman, S. L. Lee, and A. Sharma. Approximation by Lambdasplines on the circle. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):1085–1111, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ruckle:1985:STS

[524] William H. Ruckle. The strong varphi topology on symmetric sequence spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37 (??):1112–1133, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Handelman:1985:CRM

[525] David Handelman. Centres of rankmetric completions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):1134–1148, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Huneke:1985:NMC

[526] Craig Huneke and Matthew Miller. A note on the multiplicity of Cohen– Macaulay algebras with pure resolutions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):1149–1162, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ebert:1985:PPG

[527] Gary L. Ebert. Partitioning projective geometries into caps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):1163–1175, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marshall:1985:ESW

[528] Bernard Marshall. Estimates for solutions of wave equations with vanishing curvature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):1176–1200, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goulden:1985:DGJ

[529] I. P. Goulden. Directed graphs and the Jacobi–Trudi identity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):1201– 1210, ???? 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hile:1985:GST

[530] G. N. Hile and W. E. Pfaffenberger. Generalized spectral theory in complex Banach algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):1211–1236, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arthur:1985:MUV

[531] James Arthur. A measure on the unipotent variety. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(6):1237– 1274, December 1985. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vandeVel:1986:MFD

[532] M. van de Vel. Metrizability of finite-dimensional spaces with a binary convexity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1–18, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Almeida:1986:SOP

[533] Jorge Almeida. Some order properties of the lattice of varieties of commutative semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):19–47, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Larson:1986:CCR

[534] Suzanne Larson. Convexity conditions on f-rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):48–64, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kalton:1986:BEN

[535] N. J. Kalton. Banach envelopes of nonlocally convex spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):65–86, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Izuchi:1986:ABA

[536] Keiji Izuchi and Yuko Izuchi. Algebras of bounded analytic functions containing the disk algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):87–108, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lazar:1986:DHS

[537] A. J. Lazar, S.-K. Tsui, and S. Wright.

Derivations from hereditary subalgebras of C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):109–126, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bekolle:1986:PEF

[538] David Békollé. Projections sur des espaces de fonctions holomorphes dans des domaines plans. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):127–157, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roberts:1986:GCP

[539] Paul Roberts. Graded complexes over power series rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):158–178, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arthur:1986:FDO

[540] James Arthur. On a family of distributions obtained from orbits. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??): 179–214, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Srivastava:1986:SIR

[541] H. M. Srivastava and V. K. Jain. qseries identities and reducibility of basic double hypergeometric functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??): 215–231, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ng:1986:FFM

[542] Tze Beng Ng. Frame fields on manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):232–256, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berman:1986:EAM

[543] J. Berman, G. Grätzer, and C. R. Platt. Extending algebras to model congruence schemes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):257–276, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leitch:1986:NOS

[544] R. D. Leitch. A note on orbits of subgroups of the permutation groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38 (??):277–285, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brungs:1986:BDR

[545] H.-H. Brungs. Bezout domains and rings with a distributive lattice of right ideals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):286–303, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gobel:1986:SPF

[546] R. Göbel and R. Vergohsen. Solution of a problem of L. Fuchs concerning finite intersections of pure subgroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??): 304–327, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marshall:1986:FTS

[547] Bernard Marshall. The Fourier transforms of smooth measures on hypersurfaces of R^{n+1}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):328–359, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gautrin:1986:CST

[548] Henri-François Gautrin and Tapio Klemola. Continuity of the scattering transformation for the Korteweg– de Vries equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):360–375, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Radjavi:1986:TCE

[549] Heydar Radjavi. A trace condition equivalent to simultaneous triangularizability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):376–386, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chang:1986:GCG

[550] Bomshik Chang. Generators of Chevalley groups over Z. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):387–396, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Charris:1986:SOP

[551] Jairo Charris and Mourad E. H. Ismail. On sieved orthogonal polynomials. II. random walk polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??): 397–415, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maeda:1986:II

[552] Sadahiro Maeda. Isotropic immersions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38 (??):416–430, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1986:PAB

[553] J. M. Borwein and H. M. Strójwas. Proximal analysis and boundaries of closed sets in Banach space. I. theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):

431–452, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1986:CAS

[554] David Borwein and Brian Thorpe. On Cesàro and Abel summability factors for integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):453–477, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dunkl:1986:BVP

[555] Charles F. Dunkl. Boundary value problems for harmonic functions on the Heisenberg group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):478–512, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nivat:1986:ERM

[556] Maurice Nivat and Dominique Perrin.

Ensembles reconnaissables de mots biinfinis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(3):513–537, June 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Magill:1986:E

[557] K. D. Magill, Jr., P. R. Misra, and U. B.

Tewari. Epimorphisms from S(X) onto S(Y). Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):538–551, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1986:SPN

[558] Johnny A. Johnson and Gerald R.

Sherette. Structural properties of a new class of CM-lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):552–562, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Toledo:1986:SCA

[559] Juan A. Toledo. Small compact actions on chainable continua. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):563–575, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isaacs:1986:IRS

[560] I. M. Isaacs. Induction and restriction of pi-special characters. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):576–604, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahman:1986:IRC

[561] Mizan Rahman. An integral representation of a _{10}φ_{9} and continuous bi-orthogonal _{10}φ_{9} rational functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38 (??):605–618, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ginsburg:1986:AFS

[562] J. Ginsburg, I. Rival, and B. Sands. Antichains and finite sets that meet all maximal chains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):619–632, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goodearl:1986:TPD

[563] K. R. Goodearl and D. E. Handelman.

Tensor products of dimension groups and Ksb0 of unit-regular rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??): 633–658, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shelah:1986:NSB

[564] Saharon Shelah and Stevo Todorcević.

A note on small Baire spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??): 659–665, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gaveau:1986:DPH

[565] Bernard Gaveau and Jacques Vauthier.

The Dirichlet problem on the Heisenberg group. III. harmonic measure of a certain half-space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):666–671, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sherk:1986:G

[566] F. A. Sherk. The geometry of GF(q^{3}).

Canadian Journal of Mathematics = Journal canadien de mathématiques, 38 (??):672–696, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Geluk:1986:TTE

[567] J. L. Geluk, L. de Haan, and U. Stadtmüller. A Tauberian theorem of exponential type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):697–718, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Balogh:1986:PLL

[568] Zoltán T. Balogh. Paracompactness in locally Lindelöf spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):719–727, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanDulst:1986:KPN

[569] D. van Dulst and V. de Valk. (KK)- properties, normal structure and fixed points of nonexpansive mappings in Orlicz sequence spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):728–750, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Putcha:1986:SAL

[570] Mohan S. Putcha. A semigroup approach to linear algebraic groups. III. buildings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):751–768, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Magill:1986:MSL

[571] K. D. Magill, Jr. Monomorphisms of semigroups of local dendrites. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??): 769–780, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hambleton:1986:AFG

[572] Ian Hambleton and Ib Madsen. Actions of finite groups on R^{n+k} with fixed set R^{k}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):781–860, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Raphael:1986:RMI

[573] Louise A. Raphael. Resolvent means and inverting generalized Fourier transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38 (??):861–877, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

MacCluer:1986:ADC

[574] Barbara D. MacCluer and Joel H.

Shapiro. Angular derivatives and compact composition operators on the Hardy and Bergman spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??): 878–906, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pierre:1986:EPC

[575] R. Pierre. Extremal properties of constrained Tchebychev polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38 (??):907–924, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCurley:1986:SPV

[576] Kevin S. McCurley. The smallest prime value of x^{n} + a. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):925–936, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Assani:1986:OPB

[577] I. Assani. Sur les opérateurs à puissances bornées et le théorème ergodique ponctuel dans L^{p}[0, 1], 1 < p < +∞. (French) [On power-bounded operators and the timely ergodic theorem in L^{p}[0, 1], 1 < p < +∞]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):937–946, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bisztriczky:1986:SPC

[578] Tibor Bisztriczky. On the singularities of plane curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):947–968, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kumjian:1986:D

[579] Alexander Kumjian. On C^{∗}-diagonals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38 (??):969–1008, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Debiard:1986:ECR

[580] Amédée Debiard and Bernard Gaveau. Équations de Cauchy–Riemann sur SU(2) et leurs enveloppes d’holomorphie. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38 (??):1009–1024, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Buckley:1986:NEA

[581] Joseph Buckley and James Wiegold. Nilpotent extensions of abelian pgroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1025–1052, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kwasik:1986:PTS

[582] Slawomir Kwasik. On periodicity in topological surgery. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1053–1064, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lock:1986:TPO

[583] Robin H. Lock. The tensor product of operational logics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1065–1080, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marsden:1986:VTV

[584] M. J. Marsden. A Voronovskaya theorem for variation-diminishing spline approximation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1081–1093, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

May:1986:FG

[585] Coy L. May. A family of M^{∗}-groups.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 38 (??):1094–1109, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Livorni:1986:CAS

[586] Elvira Laura Livorni. Classification of algebraic surfaces with sectional genus less than or equal to six. II. ruled surfaces with dim, phisbKsbXotimesL(X) = 1. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38 (??):1110–1121, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gratzer:1986:HDL

[587] George Grätzer and Harry Lakser. Homomorphisms of distributive lattices as restrictions of congruences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??): 1122–1134, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McDonald:1986:SUS

[588] Gerard McDonald and C. Sundberg. On the spectra of unbounded subnormal operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1135–1148, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruckner:1986:PAD

[589] A. M. Bruckner, M. Laczkovich, G. Petruska, and B. S. Thomson. Porosity and approximate derivatives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??): 1149–1180, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edmunds:1986:DFO

[590] D. E. Edmunds and R. M. Edmunds. Distance functions and Orlicz– Sobolev spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1181–1198, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clarke:1986:SJC

[591] Frank H. Clarke and Vera Zeidan. Sufficiency and the Jacobi condition in the calculus of variations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1199–1209, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gervais:1986:IEF

[592] R. Gervais, Q. I. Rahman, and G. Schmeisser. (0, 2)-interpolation of entire functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1210–1227, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roberts:1986:IGC

[593] Leslie G. Roberts. The ideal generation conjecture for 28 points in P^{3}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??): 1228–1238, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See correction [626].

Akemann:1986:ESA

[594] Charles A. Akemann, Joel Anderson, and Gert K. Pedersen. Excising states of C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1239–1260, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leech:1986:FIW

[595] John Leech. Four integers whose twelve quotients sum to zero. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1261–1280, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Iliadis:1986:SWE

[596] S. Iliadis and V. Tzannes. Spaces on which every continuous map into a given space is constant. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1281–1298, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jawhari:1986:CRG

[597] El Moustafa Jawhari, Maurice Pouzet, and Ivan Rival. A classification of reflexive graphs: the use of “holes”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??): 1299–1328, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Libera:1986:BCI

[598] Richard J. Libera and Eligiusz J. Zlotkiewicz. Behavior of coefficients of inverses of α-spiral functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??): 1329–1337, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Appell:1986:TBP

[599] Jürgen Appell and Espedito De Pascale. Théorèmes de bornage pour l’opérateur de Nemyckii dans les espaces idéaux. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38 (??):1338–1355, ???? 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Belluce:1986:SAI

[600] L. P. Belluce. Semisimple algebras of infinite valued logic and bold fuzzy set theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1356–1379, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoffman:1986:HAP

[601] Peter N. Hoffman and John F.

Humphreys. Hopf algebras and projective representations of GwrSsbn and GwrAsbn. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1380–1458, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cutler:1986:HDD

[602] Colleen D. Cutler. The Hausdorff dimension distribution of finite measures in Euclidean space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1459–1484, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Apostol:1986:SPE

[603] Constantin Apostol and Lawrence Fialkow. Structural properties of elementary operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):1485–1524, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCool:1986:BCA

[604] J. McCool. On basis-conjugating automorphisms of free groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(6):1525– 1529, December 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Asmuth:1987:SSR

[605] Charles Asmuth. Some supercuspidal representations of Sp_{4}(k). Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):1–7, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ferri:1987:CSH

[606] Massimo Ferri. Colour switching and homeomorphism of manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??): 8–32, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kemp:1987:O

[607] R. R. D. Kemp. Operators on Lsb2(I)oplusC^{r}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):33–53, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bshouty:1987:VKM

[608] D. Bshouty and W. Hengartner. A variation of the Koebe mapping in a dense subset of S. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):54–73, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muhly:1987:ASN

[609] Paul S. Muhly and Kichi-Suke Saito. Analytic subalgebras of von Neumann algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):74–99, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hinton:1987:DSD

[610] D. B. Hinton and J. K. Shaw. Dirac systems with discrete spectra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??): 100–122, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

deSquire:1987:MAA

[611] Maria L. Torres de Squire. Multipliers for amalgams and the algebra Ssb0(G). Canadian Journal of Mathematics = Journal canadien de mathématiques, 39 (??):123–148, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clozel:1987:CNR

[612] Laurent Clozel. Characters of nonconnected, reductive p-adic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??): 149–167, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barr:1987:RGT

[613] Michael Barr and Michael Makkai. On representations of Grothendieck toposes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39 (??):168–221, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nicas:1987:FIH

[614] Andrew J. Nicas. Frobenius induction for higher Whitehead groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):

222–238, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lawniczak:1987:HLO

[615] Anna T. Lawniczak. High level occupation times for Gaussian stochastic processes with sample paths in Orlicz spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):239–256, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Behncke:1987:LDS

[616] Horst Behncke and George A. Elliott. Lexicographic direct sums of elementary C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):257–296, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lyzzaik:1987:MCG

[617] Abdallah Lyzzaik. The multivalent class of geometrically close-to-convex functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):297–308, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dales:1987:NIR

[618] H. G. Dales and J. P. McClure. Nonstandard ideals in radical convolution algebras on a half-line. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):309–321, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fenn:1987:MPL

[619] Roger Fenn and Denis Sjerve. Massey products and lower central series of free groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):322–337, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Friedlander:1987:DS

[620] J. B. Friedlander and H. Iwaniec.

On the distribution of the sequence n^{2}heta; (mod, 1). Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):338–344, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ranicki:1987:TTQ

[621] Andrew Ranicki. The L-theory of twisted quadratic extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??): 345–364, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sheu:1987:CTM

[622] Albert Jeu-Liang Sheu. A cancellation theorem for modules over the group C^{∗}- algebras of certain nilpotent Lie groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39 (??):365–427, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1987:PAB

[623] J. M. Borwein and H. M. Strójwas. Proximal analysis and boundaries of closed sets in Banach space. II. applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39 (??):428–472, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alpern:1987:EBE

[624] S. Alpern and V. Prasad. End behaviour and ergodicity for homeomorphisms of manifolds with finitely many ends. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):473–491, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Price:1987:STF

[625] Geoffrey L. Price. Shifts on type II_{1} factors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):492–511, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roberts:1987:CIG

[626] Leslie G. Roberts. Correction to: “The ideal generation conjecture for 28 points in P^{3},”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):512, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [593].

Tang:1987:NRC

[627] Betty Tang, Solomon W. Golomb, and Ronald L. Graham. A new result on comma-free codes of even word-length. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39 (??):513–526, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Behera:1987:PAC

[628] A. Behera and S. Nanda. mod-C Postnikov approximation of a 1connected space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):527–543, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hell:1987:ARV

[629] Pavol Hell and Ivan Rival. Absolute retracts and varieties of reflexive graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39 (??):544–567, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fricke:1987:MSG

[630] G. H. Fricke and J. A. Fridy. Matrix summability of geometrically dominated series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):568–582, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nkashama:1987:EUM

[631] M. N. Nkashama. Existence, unicité et multiplicité de solutions périodiques d’équations différentielles de Duffing non-linéaires avec dissipation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??): 583–602, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mukherjee:1987:ICM

[632] N. P. Mukherjee and Prabir Bhattacharya. On the intersection of a class of maximal subgroups of a finite group.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 39 (??):603–611, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Paterson:1987:CRT

[633] Alan L. T. Paterson. Contractive representation theory for the unitary group of C(X, M_{2}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):612–624, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hile:1987:ICB

[634] G. N. Hile and W. E. Pfaffenberger. Idempotents in complex Banach algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):625–630, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boos:1987:ITF

[635] Johann Boos. Inclusion theorems for FK-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):631–645, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hildebrand:1987:MFS

[636] Adolf Hildebrand. Multiplicative functions in short intervals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):646–672, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Meir:1987:MMB

[637] A. Meir and J. W. Moon. On major and minor branches of rooted trees. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??): 673–693, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ray:1987:RBM

[638] Gary Alan Ray. Relations between Mahler’s measure and values of L-series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39 (??):694–732, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jardine:1987:SHT

[639] J. F. Jardine. Stable homotopy theory of simplicial presheaves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):733–747, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lalonde:1987:HSD

[640] François Lalonde. Homologies de Shih:

définition et propriétés. (French) [Shih homologies: definition and properties]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39 (??):748–768, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Todorov:1987:EFC

[641] Pavel G. Todorov. Explicit formulas for the coefficients of α-convex functions, α ≥ 0. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):769–783, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kallaher:1987:MGQ

[642] Michael J. Kallaher. The multiplicative groups of quasifields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):784–793, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdahl:1987:ES

[643] R. M. Erdahl and S. S. Ryshkov. The empty sphere. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):794–824, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brownawell:1987:FPD

[644] W. Dale Brownawell. On the factorization of partial differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39 (??):825–834, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Snaith:1987:DTH

[645] Victor Snaith. A descent theorem for Hermitian K-theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):835–847, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

OMeara:1987:STS

[646] K. C. O’Meara. Simple type III selfinjective rings and rings of columnfinite matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):848–879, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bercovici:1987:TTP

[647] Hari Bercovici. Three test problems for quasisimilarity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):880–892, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mekler:1987:CTS

[648] Alan H. Mekler. Classification theory and stationary logic. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):893–907, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Salzmann:1987:CDP

[649] Helmut Salzmann. Compact 16dimensional projective planes with large collineation groups. IV. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):908–919, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wolf:1987:CCS

[650] Thomas R. Wolf. Character correspondences and pi-special characters in pi-separable groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):920–937, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aguade:1987:DFL

[651] J. Aguadé. Decomposable free loop spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):938–955, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rhemtulla:1987:EES

[652] A. H. Rhemtulla and J. S. Wilson. On elliptically embedded subgroups of soluble groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):956–968, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boileau:1987:UFA

[653] Michel Boileau and Erica Flapan. Uniqueness of free actions on S^{3} respecting a knot. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):969–982, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wimp:1987:EFA

[654] Jet Wimp. Explicit formulas for the associated Jacobi polynomials and some applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):983–1000, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jungnickel:1987:DSA

[655] Dieter Jungnickel. Divisible semiplanes, arcs, and relative difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??): 1001–1024, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Faith:1987:GTC

[656] Carl Faith. On the Galois theory of commutative rings. II. automorphisms induced in residue rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):1025– 1037, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Meyerhoff:1987:LBV

[657] Robert Meyerhoff. A lower bound for the volume of hyperbolic 3-manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??): 1038–1056, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Owa:1987:USG

[658] Shigeyoshi Owa and H. M. Srivastava. Univalent and starlike generalized hypergeometric functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??): 1057–1077, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Farnsteiner:1987:DSD

[659] Rolf Farnsteiner. Dual space derivations and H^{2}(L, F) of modular Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):1078–1106, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Murty:1987:SE

[660] M. Ram Murty and N. Saradha. On the sieve of Eratosthenes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):1107–1122, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ricker:1987:UCA

[661] Werner Ricker. Uniformly closed algebras generated by Boolean algebras of projections in locally convex spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39 (??):1123–1146, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boivin:1987:TCL

[662] Daniel Boivin. Théorèmes de convergence locale pour les résolvantes et les processus abéliens à plusieurs paramètres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):1147–1161, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Noussair:1987:SEP

[663] Ezzat S. Noussair and Charles A. Swanson. Semilinear elliptic problems with pairs of decaying positive solutions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39 (??):1162–1173, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jardine:1987:MC

[664] J. F. Jardine. The multiple Qconstruction. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):1174–1209, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Natsume:1987:IGV

[665] Toshikazu Natsume. The C^{1}-invariance of the Godbillon–Vey map in analytical K-theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):1210–1222, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Klemes:1987:IM

[666] I. Klemes. Idempotent multipliers of H^{1}(T). Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):1223–1234, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robinson:1987:CTH

[667] Derek W. Robinson. Commutator theory on Hilbert space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):1235–1280, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Debiard:1987:ARS

[668] Amédée Debiard and Bernard Gaveau.

Analysis on root systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??): 1281–1404, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Knowles:1987:SCN

[669] G. J. Knowles. On the structure of certain nest algebra modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??): 1405–1412, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Holladay:1987:ESC

[670] Kenneth Holladay. Estimating the size of context-free tiling languages. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??): 1413–1433, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fine:1987:FEB

[671] Benjamin Fine. Fuchsian embeddings in the Bianchi groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):1434–1445, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pierce:1987:CNH

[672] Stephen Pierce and Leiba Rodman. Congruences and norms of Hermitian matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):1446–1458, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ihrig:1987:HTC

[673] Edwin Ihrig. Hodge theory on compact two-dimensional spacetimes and the uniqueness of g_{ij} with a specified R_{ij}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??):1459–1474, ???? 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alpern:1987:WMM

[674] Steve Alpern and Vidhu Prasad. Weak mixing manifold homeomorphisms preserving an infinite measure. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(??): 1475–1488, ???? 1987. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Abu-Muhanna:1987:HMC

[675] Yusuf Abu-Muhanna and Glenn Schober. Harmonic mappings onto convex domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(6):1489–1530, December 1987. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kudla:1988:TCG

[676] Stephen S. Kudla and John J. Millson.

Tubes, cohomology with growth conditions and an application to the theta correspondence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1–37, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sunder:1988:S

[677] V. S. Sunder. N subspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 38–54, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berman:1988:GVF

[678] Robert D. Berman. Generalized variation and functions of slow growth. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40 (??):55–85, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Powers:1988:ITS

[679] Robert T. Powers. An index theory for semigroups of ^{∗}-endomorphisms of B(H) and type II_{1} factors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):86–114, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilman:1988:IDS

[680] Jane Gilman. Inequalities in discrete subgroups of PSL(2, R). Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):115–130, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deveney:1988:HDD

[681] James K. Deveney and Nickolas Heerema. Higher derivations and distinguished subfields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):131–141, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mayorquin:1988:MH

[682] J. G. Mayorquin. The mod 2 Khomology of Ω^{3}S^{3}X. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):142–196, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Loring:1988:TAC

[683] Terry A. Loring. K-theory and asymptotically commuting matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 197–216, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kato:1988:SEW

[684] Hisao Kato. Shape equivalences of Whitney continua of curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):217–227, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1988:TTC

[685] David Borwein and Tom Markovich. A Tauberian theorem concerning boreltype and cesàro methods of summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40 (??):228–247, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miers:1988:CPM

[686] C. Robert Miers. Commutativity preserving maps of factors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):248–256, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rieffel:1988:PMH

[687] Marc A. Rieffel. Projective modules over higher-dimensional noncommutative tori. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):257–338, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Filaseta:1988:ICP

[688] Michael Filaseta. Irreducibility criteria for polynomials with nonnegative coefficients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):339–351, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mukherjee:1988:IFM

[689] N. P. Mukherjee and Prabir Bhattacharya. On the intersection of a family of maximal subgroups containing the Sylow subgroups of a finite group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 352–359, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Guerre:1988:ITE

[690] S. Guerre and Y. Raynaud. Sur les isométries de Ł^{p}(X) et le théorème ergodique vectoriel. (French) [On isometries of Ł^{p}(X) and the vector ergodic theorem]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(2):360–391, April 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilkey:1988:EIE

[691] Peter B. Gilkey. The eta invariant and equivariant SPIN^{c} bordism for spherical space form groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):392–428, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Millet:1988:FPM

[692] Annie Millet and Louis Sucheston. On fixed points and multiparameter ergodic theorems in Banach lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 429–458, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chun:1988:ICL

[693] C. S. Chun and A. R. Freedman. Inclusions for classes of lacunary sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 459–476, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gauthier:1988:AMP

[694] P. M. Gauthier, R. Grothmann, and W. Hengartner. Asymptotic maximum principles for subharmonic and plurisubharmonic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):477–486, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lawniczak:1988:LDG

[695] Anna T. Lawniczak. Large deviations for Gaussian stochastic processes with sample paths in Orlicz spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 487–501, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beals:1988:HEN

[696] Richard Beals and Nancy K. Stanton. The heat equation for the \overline{∂}- neumann problem. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):502–512, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ringseth:1988:ACF

[697] Paul F. Ringseth. On the analytic continuation of c-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):513–531, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coupet:1988:RFH

[698] Bernard Coupet. Régularité de fonctions holomorphes sur des wedges. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 532–545, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Conner:1988:ACW

[699] P. E. Conner and Noriko Yui. The additive characters of the Witt ring of an algebraic number field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):546–588, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frankowska:1988:LBT

[700] H. Frankowska and B. Kaśkosz. Linearization and boundary trajectories of nonsmooth control systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 589–609, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Akcoglu:1988:PCA

[701] M. A. Akcoglu and L. Sucheston. Pointwise convergence of alternating sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):610–632, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kempf:1988:TPF

[702] George Kempf and Linda Ness. Tensor products of fundamental representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):633–648, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frey:1988:SGT

[703] G. Frey. On the Selmer group of twists of elliptic curves with Q-rational torsion points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):649–665, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ward:1988:QTC

[704] Doug Ward. The quantificational tangent cones. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):666–694, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aczel:1988:SDE

[705] János Aczél and Detlef Gronau. Some differential equations related to iteration theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):695–717, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kim:1988:AHF

[706] Hong Oh Kim. On an F -algebra of holomorphic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):718–741, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alpern:1988:CIM

[707] S. Alpern, J. R. Choksi, and V. S.

Prasad. Conjugates of infinite measure preserving transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 742–749, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dyn:1988:CIT

[708] N. Dyn and D. S. Lubinsky. Convergence of interpolation to transforms of totally positive kernels. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):750–768, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frohliger:1988:MHL

[709] John A. Frohliger. Maximal homotopy Lie subgroups of maximal rank. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 769–787, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sarkozy:1988:DSI

[710] A. Sárközy and C. L. Stewart. On divisors of sums of integers. IV. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 788–816, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hellegouarch:1988:NBP

[711] Yves Hellegouarch. Notions de base pour l’arithmétique de F_{q}((1/t)). (French) [Fundamentals of arithmetic in F_{q}((1/t))]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):817–832, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Packer:1988:SME

[712] Judith A. Packer. Strong Morita equivalence for Heisenberg C^{∗}-algebras and the positive cones of their K_{0}-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 833–864, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1988:SMA

[713] Lawrence G. Brown. Semicontinuity and multipliers of C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 865–988, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heywood:1988:HSR

[714] P. Heywood and P. G. Rooney. On the Hankel and some related transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):989–1009, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frappier:1988:RFI

[715] Clément Frappier. Representation formulas for integrable and entire functions of exponential type. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1010–1024, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ismail:1988:AWR

[716] Mourad E. H. Ismail and Dennis Stanton. On the Askey–Wilson and Rogers polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1025–1045, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nest:1988:CCN

[717] Ryszard Nest. Cyclic cohomology of noncommutative tori. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1046–1057, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ryshkov:1988:ESI

[718] S. S. Ryshkov and R. M. Erdahl. The empty sphere. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1058–1073, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kim:1988:WSH

[719] Hong Oh Kim and Ern Geun Kwon. Weighted subspaces of Hardy spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40 (??):1074–1083, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kulkarni:1988:CBR

[720] S. H. Kulkarni and S. Arundhathi. Choquet boundary for real function algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1084–1104, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Strichartz:1988:LAC

[721] Robert S. Strichartz. Linear algebra of curvature tensors and their covariant derivatives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1105–1143, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCool:1988:SRI

[722] J. McCool. Some remarks on IA automorphisms of free groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 1144–1155, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Furusho:1988:EPD

[723] Yasuhiro Furusho and Takaŝi Kusano.

On the existence of positive decaying entire solutions for a class of sublinear elliptic equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1156–1173, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aulaskari:1988:VRS

[724] Rauno Aulaskari. On VMOA for Riemann surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1174–1185, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fitzgerald:1988:GWR

[725] Robert W. Fitzgerald. Gorenstein Witt rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1186–1202, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahman:1988:ATS

[726] Mizan Rahman. An addition theorem and some product formulas for q-Bessel functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1203–1221, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Allegretto:1988:EPG

[727] W. Allegretto and Y. X. Huang. Existence of positive global solutions of mixed sublinear-superlinear problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40 (??):1222–1242, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beatrous:1988:PDA

[728] Frank Beatrous and Song-Ying Li. A Plessner decomposition along transverse curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1243–1255, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1988:PHS

[729] John T. Anderson. Polynomial hulls of sets invariant under an action of the special unitary group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1256–1271, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nakazi:1988:CIS

[730] Takahiko Nakazi. Certain invariant subspaces of H^{2} and L^{2} on a bidisc. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 1272–1280, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kusano:1988:RES

[731] Takaŝi Kusano, Manabu Naito, and Charles A. Swanson. Radial entire solutions of even order semilinear elliptic equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1281–1300, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ladas:1988:OSO

[732] G. Ladas, E. C. Partheniadis, and Y. G.

Sficas. Oscillations of second order neutral equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1301–1314, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hayden:1988:EAC

[733] John L. Hayden. Elementary abelian Cartesian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1315–1321, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wu:1988:PNO

[734] Pei Yuan Wu. Products of normal operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1322–1330, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kane:1988:SHC

[735] Richard Kane and Guillermo Moreno.

Spherical homology classes in the bordism of Lie groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1331–1374, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldstein:1988:UHA

[736] M. Goldstein and W. H. Ow. Uniform harmonic approximation with continuous extension to the boundary. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 1375–1388, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Felix:1988:MTL

[737] Yves Félix and Jean-Michel Lemaire. On the mapping theorem for Lusternik– Schnirel’mann category. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??): 1389–1398, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1988:PRS

[738] David F. Anderson and David E. Dobbs. Pairs of rings with the same prime ideals. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40 (??):1399–1409, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sands:1988:GHS

[739] A. D. Sands and P. N. Stewart. General heredity and strength for radical classes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40 (??):1410–1421, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baginski:1988:MGA

[740] C. Bagiński and A. Caranti. The modular group algebras of p-groups of maximal class. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1422–1435, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herrero:1988:ESQ

[741] Domingo A. Herrero. On the essential spectra of quasisimilar operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):

1436–1457, ???? 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wawrzynczyk:1988:SAU

[742] Antoni Wawrzyńczyk. Spectral analysis on upper light cone in R^{3} and the Radon transform. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):1458–1481, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wassermann:1988:EAC

[743] Antony Wassermann. Ergodic actions of compact groups on operator algebras. II. classification of full multiplicity ergodic actions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(6):1482– 1527, December 1988. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vigneras:1989:MGV

[744] Marie-France Vignéras. Moyennes galoisiennes des valeurs de fonctions L. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41 (??):1–13, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chacron:1989:VNO

[745] M. Chacron. C-valuations and normal C-orderings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):14–67, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isaacs:1989:CGA

[746] I. M. Isaacs. Coprime group actions fixing all nonlinear irreducible characters. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):68–82, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ramsamujh:1989:CND

[747] T. I. Ramsamujh. The complexity of nowhere differentiable continuous functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):83–105, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mate:1989:EFB

[748] Attila Máté and Paul Nevai. Eigenvalues of finite band-width Hilbert space operators and their application to orthogonal polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):106–122, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lustig:1989:AFG

[749] Martin Lustig, Yoav Moriah, and Gerhard Rosenberger. Automorphisms of Fuchsian groups and their lifts to free groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):123–131, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Llavona:1989:HBA

[750] J. G. Llavona and J. A. Jaramillo. Homomorphisms between algebras of continuous functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):132–162, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilligan:1989:HCM

[751] B. Gilligan, K. Oeljeklaus, and W. Richthofer. Homogeneous complex manifolds with more than one end. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41 (??):163–177, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isaacs:1989:FDI

[752] Godfrey L. Isaacs. Fractional derivatives as inverses. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):178–192, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Treger:1989:PCN

[753] Robert Treger. Plane curves with nodes.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 41 (??):193–212, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Porter:1989:PPI

[754] Jack R. Porter and R. Grant Woods.

The poset of perfect irreducible images of a space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):213–233, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schechter:1989:MO

[755] Martin Schechter. Multiplication operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):234–249, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Przytycki:1989:ECM

[756] Józef H. Przytycki. Equivalence of cables of mutants of knots. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):250–273, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Betancor:1989:MPE

[757] J. J. Betancor. A mixed Parseval’s equation and a generalized Hankel transformation of distributions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):274–284, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Casselman:1989:ISS

[758] W. Casselman. Introduction to the Schwartz space of Γ\G. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):285–320, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Philos:1989:OCS

[759] Ch. G. Philos. Oscillation criteria for second order superlinear differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):321–340, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bennett:1989:QIM

[760] F. E. Bennett. Quasigroup identities and Mendelsohn designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):341–368, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blyth:1989:RPF

[761] Russell D. Blyth and Akbar H. Rhemtulla. Rewritable products in FCby-finite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):369–384, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Casselman:1989:CEH

[762] W. Casselman. Canonical extensions of Harish-Chandra modules to representations of G. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):385–438, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Petrich:1989:CCR

[763] Mario Petrich. Congruences on completely regular semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??): 439–461, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lanteri:1989:TPC

[764] Antonio Lanteri and Daniele C. Struppa.

Topological properties of cyclic coverings branched along an ample divisor. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41 (??):462–479, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Langlands:1989:OIF

[765] R. P. Langlands and D. Shelstad. Orbital integrals on forms of SL(3). II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41 (??):480–507, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Preiss:1989:ASI

[766] D. Preiss and B. S. Thomson. The approximate symmetric integral. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??): 508–555, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roesler:1989:GM

[767] Friedrich Roesler. Generalized matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):556–576, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Melancon:1989:LWF

[768] Guy Melançon and Christophe Reutenauer. Lyndon words, free algebras and shuffles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):577–591, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fuller:1989:RB

[769] K. R. Fuller, W. K. Nicholson, and J. F. Watters. Reflexive bimodules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??): 592–611, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Landau-Treisner:1989:WCC

[770] Laurellen Landau-Treisner and Albert E. Livingston. Weakly close-to-convex meromorphic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??): 612–625, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kim:1989:RGA

[771] Yangkon Kim. Regular germs for p-adic Sp(4). Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):626–641, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fournier:1989:SUF

[772] Richard Fournier. Starlike univalent functions bounded on the real axis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41 (??):642–658, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaminska:1989:RKS

[773] A. Kamińska and B. Turett. Rotundity in Köthe spaces of vector-valued functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):659–675, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Campbell:1989:STP

[774] H. E. A. Campbell and P. S. Selick. On the semi-tensor product of the Dyer– Lashof and Steenrod algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??): 676–701, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1989:ENP

[775] Jonathan M. Borwein and Simon Fitzpatrick. Existence of nearest points in Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):702–720, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhang:1989:SPI

[776] Shuang Zhang. On the structure of projections and ideals of corona algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):721–742, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Krattenthaler:1989:ATV

[777] Christian Krattenthaler. q-analogue of a two variable inverse pair of series with applications to basic double hypergeometric series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):743–768, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohn:1989:BIC

[778] William S. Cohn. Boundary interpolation for continuous holomorphic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):769–785, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Giordano:1989:RAA

[779] T. Giordano and D. E. Handelman.

Real AF C^{∗}-algebras with K_{0} of small rank. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):786–807, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Powers:1989:FSS

[780] Victoria Powers. Finite spaces of signatures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):808–829, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Banaschewski:1989:CPM

[781] B. Banaschewski and A. Pultr. Cauchy points of metric locales. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):830–854, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gora:1989:CID

[782] P. Góra and A. Boyarsky. Compactness of invariant densities for families of expanding, piecewise monotonic transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(5):855–869, October 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Choa:1989:CNB

[783] Jun Soo Choa and Hong Oh Kim.

Composition with a nonhomogeneous bounded holomorphic function on the ball. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):870–881, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haagerup:1989:FAF

[784] Uffe Haagerup and Gilles Pisier. Factorization of analytic functions with values in noncommutative L_{1}-spaces and applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):882–906, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dostal:1989:SPM

[785] Milos Dostál and Bernard Gaveau. The stationary phase method for certain degenerate critical points. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??): 907–931, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Conner:1989:R

[786] P. E. Conner and Jürgen Hurrelbrink. The 4-rank of K_{2}(O). Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):932–960, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roddy:1989:WPO

[787] Michael S. Roddy. On the word problem for orthocomplemented modular lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):961–1004, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Biancofiore:1989:HSB

[788] Aldo Biancofiore. On the hyperplane sections of blow-ups of complex projective plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):1005–1020, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Phillips:1989:CAT

[789] N. Christopher Phillips. Classifying algebras for the K-theory of σ-C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):1021–1089, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aupetit:1989:SAJ

[790] Bernard Aupetit and Line Baribeau.

Sur le socle dans les algèbres de Jordan–Banach. (French) [On the basis in Jordan–Banach algebras]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??): 1090–1100, ???? 1989. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kuo:1989:GNP

[791] Tzee Char Kuo. Generalized Newton– Puiseux theory and Hensel’s lemma in C[![x, y]!]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):1101–1116, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lam:1989:NFP

[792] C. W. H. Lam, L. Thiel, and S. Swiercz.

The nonexistence of finite projective planes of order 10. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):1117–1123, ???? 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ruttimann:1989:AJH

[793] Gottfried T. Rüttimann. The approximate Jordan–Hahn decomposition. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41 (6):1124–1146, December 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).